### v2

# Explainable Molecular Property Prediction: Aligning Chemical Concepts with Predictions via Language Models

Abstract
Providing explainable molecular property predictions is critical for many scientific domains, such as drug discovery and material science. Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships. In this work, we develop a new framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole.. We first leverage a designated molecular representation — the Group SELFIES — as it can provide chemically meaningful semantics. Because attention mechanisms in Transformers can inherently capture relationships within the input, we further incorporate the attention weights and gradients together to generate explanations for capturing the functional group interactions. We then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations. We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations. Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.8\%$, being the state-of-the-art in explainable molecular property prediction.
1 Introduction
Molecular property prediction aims to reveal the relationships between molecular structures and their properties, assisting scientists in screening molecules for various applications such as drug discovery and material design [1, 2, 3, 4, 5]. Several learning-based models are devised based on the underlying molecular representations, such as graph-based molecular representations and string-based molecular representations. Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES [6]), stand out for their simplicity and adaptability [2, 7, 8]. By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert [9] offer higher throughput and accuracy for molecular property prediction [2, 10].
Figure 1: (a) The molecule visualization of prediction/explanation. The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule. (b)-(e) are the explanation results obtained with various methods: (b) OrphicX [11]; (c) GNNExplainer [12], (d) GNN with gradient-based explainability technique (GradCAM [13]); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES [8] as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains. Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions. These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design [14, 15, 16]. With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction [17, 18, 19, 20].
Figure 2: The SMILES and Group SELFIES strings of p-nitrobenzoic acid molecule ($\mathrm{C_{7}H_{5}NO_{4}}$): The tokens in the Group SELFIES string highlighted by color are the corresponding functional groups.
However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously. Specifically, 1) from the molecular representation perspective, the commonly used representations do not explicitly encode the chemically meaningful substructures; current explainability methods can only highlight individual atoms and bonds as explanations (see Fig. 1 (b)$\sim$(e)). 2) From the perspective of explainability techniques, current methods suffer from two main limitations. First, they cannot effectively capture the interactions between functional groups within the molecular structure. Second, they could not generate explanations that align with the domain knowledge of chemists. As a result, they fail to produce explanations that faithfully reflect the structure-property relationships. (see Fig. 1 (f)). Therefore, a novel framework is imperative to provide plausible explanations for molecule property predictions.
The recently proposed string-based molecular representation — Group SELFIES [8] — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations. As shown in Fig. 2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes functional groups as tokens, including a benzene, a nitro group, and a carboxyl group. Therefore, with the rich chemical semantics of Group SELFIES, this work develops a new explainable molecular property prediction framework that can provide chemical concepts-aligned explanations based on language models, called Lamole (see Fig. 1 (g)). In particular, we address the above issues with the following contributions.
  1. 1.
    Lamole uses the molecular representation — Group SELFIES — as input tokens for providing the chemically meaningful semantics. By deriving the gradients regarding the input functional group/fragment tokens, gradient-based explanations are generated to indicate the contribution of each functional group/fragment to the molecular properties.
  2. 2.
    By disentangling the information flows of Transformer-based LMs, we observe that the gradient-based explanations cannot well capture the functional group interactions. Therefore, we further integrate attention weights into gradients to generate explanations that can perceive the functional group interactions. The intuition is that the attention weights in the information flows can inherently capture the functional group interactions within the input.
  3. 3.
    To make the explanations more faithfully respect the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations. We show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%.
  4. 4.
    We first bridge the manifold hypothesis with explainable molecular property prediction. We theoretically demonstrate that the elaborated marginal loss aligns explanations with the data manifold, respecting the structure-property relationship.
Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.8\%$. We also quantitatively evaluate explanations based on the first proposed plausibility metric. Compared to alternative baselines, the explanation plausibility of Lamole is improved by up to $9\%$. Extensive experimental studies demonstrate Lamole achieves state-of-the-art performance in explainable molecular property prediction.
2 Related Work
Explainable Molecular Property Prediction. To gain insights into the model predictions, several explainable GNNs are proposed to explain the relationship between the input molecular graph and the prediction [11, 12, 21, 20]. However, as shown in Fig. 1 (b)$\sim$(d), they only provide explanations at individual atoms and bond levels instead of chemically meaningful substructures. On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction [10, 22, 23, 24]. However, the "black-box" characteristics of LMs hamper trust and efficient use of these potent computational tools in scientific domains.
Explainability Techniques in LMs. Transformers, lying the cornerstone of LMs, have advanced the state-of-the-art in various research fields, including molecular property prediction [10, 22, 24]. One way to explain their output is to use the attention weights over the input tokens. However, recent studies suggest that "attention is not explanation" because attention weights are often inconsistent with the feature importance [25, 26, 27]. Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the relationship between the input and output. However, the generated explanations may change drastically with very small perturbations [28]. Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature [13]. However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations [29, 28, 30]. To alleviate this issue, by disentangling the model architectures’ information flows, we combine the attention weights and gradients to generate explanations that faithfully reveal the structure-property relationship.
3 Methodology
Problem Setup. Given a molecular graph $g^{(i)}\in\mathcal{G}=\{(g^{(i)},y^{(i)})\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as $f:g\rightarrow y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$. Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig. 2. In addition, this work calibrates the explanations in a supervised manner. For this purpose, a few molecules with annotation masks are provided. The annotation masks $\mathbf{m}(g^{(i)})\in\{0,1\}^{j}$ indicate whether $t_{j}^{(i)}$ is the token of a ground truth substructure, where $\mathbf{m}_{j}(g^{(i)})=1$ denotes the substructure corresponding to the token $t_{j}^{(i)}$ inducing the molecular property of $g^{(i)}$. This work uses $\mathcal{D}=\{(s^{(i)},y^{(i)})\}$ with a few annotation masks to learn a model $f$ for explainable molecular property prediction. We omit the superscript (i) for simplicity in the following parts.
Figure 3: An illustration of Lamole. Left panel: Group SELFIES strings are tokenized for fine-tuning the pre-trained language model, and an MLP classifier is equipped with a cross-entropy loss $\mathcal{L}_{CE}$ for molecular property prediction. Right panel: We disentangle the information flows of the Transformer to assert that both attention weights and gradient determine the output. Therefore, we incorporate the attention weights and gradients together to generate importance scores $\mathbf{v}$ as explanations. In addition, a marginal loss $\mathcal{L}_{M}$ is designed to align explanations with the chemists’ annotations $\mathbf{m}$.
3.1 Our Design: Lamole
In this work, we use pre-trained Transformer-based LMs, e.g., Bert family models, as our backbone. Then, we fine-tune the LMs on the Group SELFIES corpus for explainable molecular property prediction. An illustration of proposed Lamole is in Fig. 3. In what follows, we will introduce the detailed design of Lamole.
Fine-Tuning Stage. Fig. 3 shows the fine-tuning stage of the proposed Lamole. We assume that the Transformer encoder in Lamole stacks $L$ identical Transformer layers to encode the molecular string $s$ as token embedding $\mathbf{h}^{(l)}=\{\mathbf{h}_{1}^{(l)},\ldots,\mathbf{h}_{j}^{(l)}\}$, where $\mathbf{h}^{(l)}$ is the token embedding at the $l$-th layer. We use the self-attention weighted average embedding $\mathbf{h}_{o}=(\sum_{j=1}\alpha_{j}\cdot\mathbf{h}_{j}^{(L)})/(\sum_{j=1}\alpha_{j})$ for molecular property prediction, where $\alpha_{j}$ is the attention weight of the $j$-th token. A multilayer perceptron (MLP) classifier is added to predict the molecular property $\hat{y}=\mathrm{MLP}(\mathbf{h}_{o})$ by minimizing the classification cross-entropy loss $\mathcal{L}_{CE}(y,\hat{y})$.
Information Flow-Based Explanations. To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature [13]. However, we argue that using gradients alone cannot effectively capture functional group interactions. As depicted in Fig. 1 (f), the gradient-based method CradCAM incorrectly attributes the property to the nitro group and a benzene ring that is not connected to the nitro group.
We illustrate the possible reason by disentangling the information flows of the Transformer. As shown in Fig. 3 right panel, due to the skip connection and attention mechanism, both attention weights, gradients, and the input contribute to the outputs. Therefore, using gradients alone as explanations could fall short of capturing interactions. To address the issue, we leverage both attention and gradients, as well as the input, to derive explanations. Below, we will elaborate on integrating attention weights into gradient-based explanations.
Firstly, we show the process of deriving the gradient-based explanations. Similar to GradCAM [13], the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by
$\nabla\mathbf{h}_{j}^{(l)}=\frac{\partial\hat{y}}{\partial\mathbf{h}_{j}^{(l)}},$ (1)
where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$. Due to the skip connection in Fig. 3 right panel, the input and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
$\mathbf{w}_{j}^{(l)}=\nabla\mathbf{h}_{j}^{(l)}\odot\mathbf{h}_{j}^{(l)},$ (2)
where $\odot$ is the Hadamard product. The weighted importance is regarded as the gradient-based explanation.
The interaction among tokens can be revealed by the self-attention mechanism in Fig. 3 right panel. The attention mechanism calculates pairwise similarity scores between all pairs of tokens to determine attention weights, and these attention weights inherently encode the functional group interactions. Therefore, we combine the attention weights with the gradient-based explanation to capture functional group interactions. Assuming the attention weights of the $j$-th token at the $l$-th layer is $\alpha_{j}^{(l)}$, we integrate the attention weights with weighted importance $\mathbf{w}$ to consider the interactions. The importance score of the molecule $g$ can be obtained by
$\mathbf{v}_{j}^{(l)}(g)=\left(\tanh(\overline{\alpha_{j}^{(l)}})\cdot\tanh(\mathbf{w}_{j}^{(l)})\right)^{\frac{1}{2}},$ (3)
where $\overline{\alpha_{j}^{(l)}}$ is the averaging of attention weights of multiple attention heads. Finally, we sum $\mathbf{v}_{i}^{(l)}$ over all layers as the final importance score of the $j$-th token,
$\mathbf{v}(g)=\mathrm{softmax}(\sum_{l=1}^{L}\mathbf{v}^{(l)}(g)).$ (4)
The higher the importance scores, the greater the contribution of the corresponding functional groups/fragments to the molecular property.
Towards Plausible Explanations. One plausible explanation should faithfully uncover the structure-property relationships. In other words, the explanation should match the ground-truth substructures with high confidence. Nevertheless, the importance scores of ground-truth substructures might not be significantly higher than those of other parts. To address this issue, we propose a marginal loss to explicitly align explanations with the chemists’ annotations to improve the explanations’ plausibility.
First, we formally define the plausibility of explanations. "Plausibility" refers to how the interpretation convinces humans [31, 32, 33]. Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
Definition 1
(Plausibility): Given the importance scores $\mathbf{v}$ over all tokens in the molecule $g$, the mean importance score $\mathbf{v}_{\in\mathcal{T}_{g}}$ over ground truth substructures $\mathcal{T}_{g}$ and the mean importance score $\mathbf{v}_{\notin\mathcal{T}_{g}}$ over other substructures $\mathcal{T}_{g}$ are denoted by $\mathbf{v}_{\in\mathcal{T}_{g}}=\frac{\sum_{j}\mathbf{v}_{j}\cdot\mathbb{I}(t_{j}\in\mathcal{T}_{g})}{\sum_{j}\mathbb{I}(t_{j}\in\mathcal{T}_{g})}$ and $\mathbf{v}_{\notin\mathcal{T}_{g}}=\frac{\sum_{j}\mathbf{v}_{j}\cdot\mathbb{I}(t_{j}\notin\mathcal{T}_{g})}{\sum_{j}\mathbb{I}(t_{j}\notin\mathcal{T}_{g})}$, respectively, where $\mathbb{I}(\cdot)$ is the indicator function. The explanations’ plausibility $\mathrm{EP}(g)$ is defined as the ratio of the difference between $\mathbf{v}_{\in\mathcal{T}_{g}}$ and $\mathbf{v}_{\notin\mathcal{T}_{g}}$ to $\mathbf{v}_{\notin\mathcal{T}_{g}}$, $\mathrm{EP}(g)=\frac{\mathbf{v}_{\in\mathcal{T}_{g}}-\mathbf{v}_{\notin\mathcal{T}_{g}}}{\mathbf{v}_{\notin\mathcal{T}_{g}}}.$ (5)
The higher the $\mathrm{EP}$ value, the greater the confidence of the explanation in matching the ground truth substructure.
In our work, the ground truth substructures are annotated by a binary mask vector $\mathbf{m}(g)\in\{0,1\}^{j}$. It is worth noting that using only a few annotations can significantly improve the explanation accuracy. Specifically, the mask vector $\mathbf{m}$ enforces the explanations to align with the ground truth substructures. To achieve the goal, a max-margin loss is designed by maximizing the mean value of the importance scores of tokens that have mask values of 1 while minimizing the mean value of importance scores for tokens with mask values of 0.
$\mathcal{L}_{M}(\mathbf{v},\mathbf{m})=\mathbb{E}_{g\in\mathcal{G}}\left[\max\left(0,\frac{\sum_{j=1}(1-\mathbf{m}_{j}(g))\cdot\mathbf{v}_{j}(g)}{N_{s}}-\frac{\sum_{j=1}\mathbf{m}_{j}(g)\cdot\mathbf{v}_{j}(g)}{N_{c}}\right)+\triangle_{1}\right],$ (6)
where $\triangle_{1}$ is a margin term, $N_{s}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 0, and $N_{c}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 1. The overall optimization objective of the fine-tuning stage is to minimize $\mathcal{L}_{CE}+\mathcal{L}_{M}$. By minimizing $\mathcal{L}_{M}$, Lamole aligns the explanations with the domain knowledge, enhancing the plausibility of explanations.
The next section will theoretically show that by using the designed marginal loss, the explanations can faithfully reflect the structure-property relationships.
3.2 Theoretical Analysis
We bridge the manifold hypothesis with the marginal loss to theoretically show that the explanations can respect the structure-property relationships. Before giving the proof, the notation and definition regarding the manifold hypothesis are presented.
Manifold Hypothesis. It is widely believed that natural data, including molecules, distribute around a manifold [34, 11, 35, 36]. According to the manifold hypothesis for gradient-based explanations [34], if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained. We call these features "causal features" in our work. Conversely, if a feature is orthogonal to the manifold, then the feature does not contribute to the class. We call these features "spurious features".
With the annotation masks, the causal features $s^{*}$ and spurious features $\overline{s}^{*}$ can be distinguished by $s^{*}=s\odot\mathbf{m}(g)$ and $\overline{s}^{*}=s\odot(1-\mathbf{m}(g))$, respectively, where $\overline{s}^{*}\cup s^{*}=s$ and $s^{*}\cap\overline{s}=\varnothing$. By projecting the causal features and spurious features into the data manifold $\mathcal{M}$, the corresponding manifold regarding the causal features and spurious features can be defined as follows,
Definition 2
(Causal feature manifold and spurious feature manifold): Assume the distribution $p(g|y)$ is implicitly modeled by a manifold $\mathcal{M}$, and the manifold can be decomposed into two components, $\underbrace{p(g|y)}_{\mathcal{M}}=\underbrace{p(g|y)\odot\mathbf{m}(g)}_{\mathcal{M}_{c}}+\underbrace{p(g|y)\odot(1-\mathbf{m}(g))}_{\mathcal{M}_{s}},$ (7) where $\mathcal{M}_{c}$ is the causal feature manifold and $\mathcal{M}_{s}$ represents the spurious feature manifold.
With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships. Due to the page limitation, we provide the proof in Appendix 1.
Theorem 1
The marginal loss of Eq. (6) aligns the gradient-based explanations $\nabla_{g}\log p(y|g)$ with the tangent space of the causal feature manifold $\mathcal{M}_{c}$, thus respecting the structure-property relationships.
4 Experiments
4.1 Experimental Setup
Datasets. We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction. Six mutagenicity datasets are Mutag [37], Mutagen [38], PTC-FM [39], PTC-FR [39], PTC-MM [39], and PTC-MR [39]. For hepatotoxicity [39], the Liver dataset [40] is used. We used these datasets as the ground truth substructures in these datasets are known. For detailed information on these datasets, please refer to Appendix 2.
Evaluation Metrics. We use three metrics for evaluating the performance of the proposed Lamole.
  1. 1.
    Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$.
  2. 2.
    Explanation Accuracy: We follow the experimental settings in GNNExplainer [12], which formulates the explanation problem as a binary classification of edges. We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation. We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones.
  3. 3.
    Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
Table 1: Mean Classification Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
GCN [41] 84.6 78.9 54.8 63.0 57.8 53.3 41.1
DGCNN [42] 85.8 74.8 57.3 63.5 61.0 58.6 44.6
edGNN [43] 86.9 75.2 59.8 65.7 64.4 56.3 44.5
GIN [44] 87.5 82.3 62.1 66.2 65.1 64.0 44.9
RW-GNN [45] 87.2 80.3 61.9 64.0 62.4 57.0 43.2
DropGNN [46] 89.4 80.7 62.0 66.0 63.7 64.2 45.0
IEGN [47] 84.6 80.1 60.8 59.8 61.1 59.5 45.3
ChemBERTa [10] 86.8 78.0 60.0 65.7 60.4 58.7 45.7
Lamole+DistilBert 84.2 76.8 57.5 69.0 60.2 64.5 47.2
Lamole+DeBerta 86.8 73.7 58.6 69.5 59.7 63.8 45.8
Lamole+Bert 88.2 74.5 62.4 70.0 61.2 66.0 47.5
  • and denote the best result and second result, respectively.
Baselines. We combine Lamole into three BERT family models, such as DistilBert [48], DeBerta [49], and Bert to evaluate the performance of the proposed Lamole. For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa [10] and several GNNs including GCN [41], DGCNN [42], edGNN [43], GIN [44], RW-GNN [45], DropGNN [46], and IEGN [47].
For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad [50], CAM [51], GradInput [52], and GradCAM [13], 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX [11], GNNExplainer [12], PGExplainer [21], and Gem [20].
Implementation Details. Following ChemBERTa [10], our used BERT family models are pre-trained on 100,000 molecules from the ZINC dataset [53]. 15% tokens in each input string are randomly for masked language learning. We finetuned the models on Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, PTC-MR, and Liver. For each dataset, the ratio of samples with annotation masks over the data size is 10%. The details of experimental settings can be found in Appendix 3.
4.2 Results
Table 2: Mean Explanation Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
SmoothGrad+GCN [50] 68.0 66.1 66.8 62.2 64.5 61.9 66.0
CAM+GCN [51] 66.8 65.4 66.9 62.7 61.4 65.5 67.5
GradInput+GCN [52] 69.1 65.1 68.5 64.3 62.3 62.0 67.2
GradCAM+GCN [13] 68.3 65.7 70.4 65.4 65.8 66.6 68.4
SmoothGrad+Bert [50] 71.3 71.3 72.8 70.6 65.2 68.9 74.4
CAM+Bert [51] 70.1 69.8 66.3 69.4 63.9 66.2 72.8
GradInput+Bert [52] 73.1 71.1 72.1 66.6 64.0 68.1 75.0
GradCAM+Bert [13] 73.5 71.5 76.3 69.0 69.1 71.2 75.8
OrphicX [11] 76.4 70.4 71.1 68.4 69.0 69.1 73.0
GNNExplainer [12] 69.5 63.0 68.6 67.3 66.0 66.3 71.4
PGExplainer [21] 65.7 58.0 69.3 66.3 64.7 65.1 71.0
Gem [20] 72.8 65.5 69.8 66.8 68.2 68.3 72,5
Lamole+DistilBert 70.9 73.0 74.0 70.2 69.6 78.1 76.1
Lamole+DeBerta 76.1 75.0 79.9 72.1 70.3 77.2 75.0
Lamole+Bert 77.8 75.2 81.1 72.2 72.0 73.1 77.3
Figure 4: Explanation visualization of one molecule (ID: 155) from the Mutag dataset.
(a) Mutag
(b) PTC-FM
(c) PTC-MR
Figure 5: The explanation accuracy of Lamole with different annotation rates on the Mutag, PTC-FM, and PTC-MR datasets.
(a) Mutag
(b) PTC-FM
(c) PTC-MR
Figure 6: The explanation plausibility of the compared algorithms on the Mutag, PTC-FM, and PTC-MR datasets.
Molecular Property Prediction. Table 1 shows the classification accuracy of compared algorithms. Our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert slightly outperform most compared algorithms on the Mutag, PTC-FM, PTC-FR, PTC-MR, and Liver datasets. Though the LMs, including Lamole and ChemBERTa, demonstrate good performance for molecular property prediction, they are inferior to GNN-based methods on the Mutagen and PTC-MM datasets. In addition, Lamole models show superior performance over ChemBERTa. This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the structure-property relationships.
Explanation Performance. Table 2 presents the explanation accuracy of the compared explainability techniques. It is obvious that the LM-based methods using the Group SELFIES strings demonstrate significantly higher explanation accuracy over explainable GNNs. Specifically, Lamole improves the explanation accuracy by $1.4\%\sim 14.8\%$ compared to the baseline methods. This is because the molecular graph-based explainable GNNs can only highlight individual atoms and bonds rather than chemically meaningful substructures like functional groups. Instead, Group SELFIES explicitly encode the chemically meaningful substructures as tokens, and using Group SELFIES can help explainable methods to generate chemically meaningful explanations by highlighting the corresponding tokens.
We also selected some representative molecules for explanation visualization. These explanations are shown in Figs. 1, 4, and 12$\sim$15 (see Appendix), respectively. The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments. As shown in Fig. 4, our proposed Lamole assigns the highest two importance scores to the two ground truth function groups, i.e., benzene-1 and nitro group. In contrast, OrphicX cannot provide chemically meaningful explanations, and CradCAM+Bert cannot accurately identify the functional group interactions. Compared to baseline methods, Lamole provides chemically meaningful explanations. Particularly, the interaction among the functional groups is successfully captured. The visualization demonstrates Lamole’s superior interpretation in uncovering the structure-property relationships.
We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%), as presented in Figs. 5 and 10 (see Appendix). Clearly, using more molecules with ground truth annotations can continuously improve explanation accuracy. In addition, compared to the results without using the marginal loss $\mathcal{L}_{M}$, using only 10% molecules with ground truth annotations can significantly improve explanation accuracy by up to 5%, confirming the effectiveness of using annotations. Though increasing the annotated ground truth rate can only bring a limited increase in the explanation accuracy, there is a trade-off between the achieved explanation accuracy and the additional annotation costs.
In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$. The statistical results of $\mathrm{EP}$ are presented in Figs. 6 and 11 (see Appendix). From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property. Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$. The statistical results of $\mathrm{EP}$ in Figs. 6 and 11 (see Appendix) indicate Lamole can provide explanations with higher confidence in matching the ground truth substructures.
Due to the page limitation, more analysis regarding the ablation studies on attention weights and the marginal loss, the impact of annotation rates, the analysis regarding the explanation plausibility, and more visualization results can be found in the Appendix.
5 Conclusions
This work developed a new framework for explainable molecular property prediction based on language models, called Lamole. Lamole uses Group SELFIES as input for perceiving the chemically meaningful semantics. By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations that can capture the functional group interactions. Furthermore, one marginal loss is designed to calibrate the explanations by aligning them with the chemists’ annotation. With the above designs, Lamole can provide chemically meaningful explanations while faithfully uncovering the structure-property relationships. Lamole ’s effectiveness has been demonstrated through theoretical analysis and extensive experimental validation.

### v3

@@
 Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships.
-In this work, we develop a new framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole..
-We first leverage a designated molecular representation — the Group SELFIES — as it can provide chemically meaningful semantics.
-Because attention mechanisms in Transformers can inherently capture relationships within the input, we further incorporate the attention weights and gradients together to generate explanations for capturing the functional group interactions.
-We then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations.
+In this work, we develop a framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole..
+We take a string-based molecular representation — Group SELFIES — as input tokens to pre-train and fine-tune our Lamole, as it provides chemically meaningful semantics.
+By disentangling the information flows of Lamole, we propose considering both self-attention weights and gradients for better quantification of each chemically meaningful substructure’s impact on the model’s output.
+To make the explanations more faithfully respect the structure-property relationship, we then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations.
 We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations.
-Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.8\%$, being the state-of-the-art in explainable molecular property prediction.
+Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.3\%$, being the state-of-the-art in explainable molecular property prediction.
 1 Introduction
-Molecular property prediction aims to reveal the relationships between molecular structures and their properties, assisting scientists in screening molecules for various applications such as drug discovery and material design [1, 2, 3, 4, 5].
-Several learning-based models are devised based on the underlying molecular representations, such as graph-based molecular representations and string-based molecular representations.
-Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES [6]), stand out for their simplicity and adaptability [2, 7, 8].
-By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert [9] offer higher throughput and accuracy for molecular property prediction [2, 10].
+Molecular property prediction aims to reveal the molecular structures-property relationships, assisting scientists in screening molecules for various applications such as drug discovery and material design (Fang et al., 2022; Deng et al., 2023; Tripp et al., 2023; Wang et al., 2024; Ekström Kelvinius et al., 2023; Hong et al., 2024).
+Several learning-based models are devised based on the underlying molecular representations, such as graph-based and string-based molecular representations.
+Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES (Weininger, 1988)), stand out for their simplicity and adaptability (Deng et al., 2023; Wigh et al., 2022; Cheng et al., 2023).
+By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert (Kenton & Toutanova, 2019) offer higher throughput and accuracy for molecular property prediction (Deng et al., 2023; Chithrananda et al., 2020).
 Figure 1: (a) The molecule visualization of prediction/explanation.
 The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule.
-(b)-(e) are the explanation results obtained with various methods: (b) OrphicX [11]; (c) GNNExplainer [12], (d) GNN with gradient-based explainability technique (GradCAM [13]); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES [8] as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
+(b)-(e) are the explanation results obtained with various methods: (b) OrphicX (Lin et al., 2022); (c) GNNExplainer (Ying et al., 2019), (d) GNN with gradient-based explainability technique (GradCAM (Selvaraju et al., 2017)); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES (Cheng et al., 2023) as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
+Figure 2: The SMILES and Group SELFIES strings of p-nitrobenzoic acid molecule ($\mathrm{C_{7}H_{5}NO_{4}}$): The tokens in the Group SELFIES string highlighted by color are the corresponding functional groups.
 Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains.
 Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions.
-These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design [14, 15, 16].
-With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction [17, 18, 19, 20].
-Figure 2: The SMILES and Group SELFIES strings of p-nitrobenzoic acid molecule ($\mathrm{C_{7}H_{5}NO_{4}}$): The tokens in the Group SELFIES string highlighted by color are the corresponding functional groups.
+These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design (Wu et al., 2023; Wellawatte et al., 2023; Das et al., 2022).
+With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction (Proietti et al., 2024; Ying et al., 2019; Ye et al., 2023; Lin et al., 2021).
 However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously.
@@
 First, they cannot effectively capture the interactions between functional groups within the molecular structure.
-Second, they could not generate explanations that align with the domain knowledge of chemists.
+Second, they could not generate explanations that align with chemists’ intuition.
 As a result, they fail to produce explanations that faithfully reflect the structure-property relationships.
@@
 1 (f)).
-Therefore, a novel framework is imperative to provide plausible explanations for molecule property predictions.
-The recently proposed string-based molecular representation — Group SELFIES [8] — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations.
+Therefore, an effective framework is imperative for explainable molecule property predictions.
+The recently proposed string-based molecular representation — Group SELFIES (Cheng et al., 2023) — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations.
 As shown in Fig.
-2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes functional groups as tokens, including a benzene, a nitro group, and a carboxyl group.
-Therefore, with the rich chemical semantics of Group SELFIES, this work develops a new explainable molecular property prediction framework that can provide chemical concepts-aligned explanations based on language models, called Lamole (see Fig.
+2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes chemically meaningful substructures as tokens, including a benzene, a nitro group, and a carboxyl group.
+Compared with 2D molecular graphs, Group SELFIES provides inherent semantic information, making it easier for the model to capture and understand chemically meaningful semantics.
+Moreover, using Group SELFIES eliminates the need to identify or segment chemically meaningful substructures in 2D molecular graphs.
+With Group SELFIES’s simplicity and adaptability, this work develops an explainable molecular property prediction framework based on language models to provide chemical concepts-aligned explanations, called Lamole (see Fig.
 1 (g)).
-In particular, we address the above issues with the following contributions.
+The contributions can be summarized as follows.
   1. 1.
-    Lamole uses the molecular representation — Group SELFIES — as input tokens for providing the chemically meaningful semantics.
-By deriving the gradients regarding the input functional group/fragment tokens, gradient-based explanations are generated to indicate the contribution of each functional group/fragment to the molecular properties.
+    We found that existing explainability techniques fail to provide chemically meaningful explanations, perceive functional group interactions, and reveal molecular structure-property relationships faithfully.
+To address the issues, we use Group SELFIES to pre-train and fine-tune LMs to make LMs easily understand the chemically meaningful semantics.
+Moreover, the process of generating explanations should reflect the reasoning process behind the model architecture.
+Therefore, by disentangling the information flows of Transformer-based LMs, we integrate the self-attention weights and gradients to capture the substructure interactions to better quantify each chemically meaningful substructure’s contribution to the predicted molecular properties.
   2. 2.
-    By disentangling the information flows of Transformer-based LMs, we observe that the gradient-based explanations cannot well capture the functional group interactions.
-Therefore, we further integrate attention weights into gradients to generate explanations that can perceive the functional group interactions.
-The intuition is that the attention weights in the information flows can inherently capture the functional group interactions within the input.
-  3. 3.
     To make the explanations more faithfully respect the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations.
 We show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%.
-  4. 4.
+  3. 3.
     We first bridge the manifold hypothesis with explainable molecular property prediction.
 We theoretically demonstrate that the elaborated marginal loss aligns explanations with the data manifold, respecting the structure-property relationship.
-Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.8\%$.
+Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.3\%$.
 We also quantitatively evaluate explanations based on the first proposed plausibility metric.
@@
 2 Related Work
-Explainable Molecular Property Prediction.
-To gain insights into the model predictions, several explainable GNNs are proposed to explain the relationship between the input molecular graph and the prediction [11, 12, 21, 20].
-However, as shown in Fig.
-1 (b)$\sim$(d), they only provide explanations at individual atoms and bond levels instead of chemically meaningful substructures.
-On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction [10, 22, 23, 24].
-However, the "black-box" characteristics of LMs hamper trust and efficient use of these potent computational tools in scientific domains.
-Explainability Techniques in LMs.
-Transformers, lying the cornerstone of LMs, have advanced the state-of-the-art in various research fields, including molecular property prediction [10, 22, 24].
-One way to explain their output is to use the attention weights over the input tokens.
-However, recent studies suggest that "attention is not explanation" because attention weights are often inconsistent with the feature importance [25, 26, 27].
-Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the relationship between the input and output.
-However, the generated explanations may change drastically with very small perturbations [28].
-Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature [13].
-However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations [29, 28, 30].
-To alleviate this issue, by disentangling the model architectures’ information flows, we combine the attention weights and gradients to generate explanations that faithfully reveal the structure-property relationship.
+Several explainable GNNs are proposed to explain the relationship between the input graph and the prediction (Sun et al., 2023; Xiong et al., 2019; Lin et al., 2022; Ying et al., 2019; Luo et al., 2020; Lin et al., 2021).
+Among these works, structure similarity or attention weights are proposed to capture structural interaction (Sun et al., 2023; Xiong et al., 2019).
+However, two similar substructures do not necessarily lead to interaction between them, and attention weights are often inconsistent with the feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020).
+In addition, as shown in Fig.
+1 (b)$\sim$(d), some trivial structures received relatively high importance scores, indicating the explanations might not align well with the chemical concepts.
+On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction (Chithrananda et al., 2020; Wang et al., 2019; Ahmad et al., 2022; Ross et al., 2022).
+However, the "black-box" characteristics of LMs hamper trust use of these potent computational tools in scientific domains.
+Some explainability techniques could be applied to LMs.
+One way is to use the attention weights over the input tokens.
+However, recent studies suggest that "attention is not explanation" because attention weights could not reflect the true feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020).
+Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the input importance.
+However, the generated explanations may change drastically with very small perturbations (Agarwal et al., 2021).
+Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature (Selvaraju et al., 2017).
+However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations (Adebayo et al., 2018; Agarwal et al., 2021; Rudin, 2019).
+Therefore, the explanation generation process should reflect the model reasoning process behind the model architectures.
+To this end, we disentangle the model architectures’ information flows to generate explanations that faithfully reveal the structure-property relationship.
 3 Methodology
 Problem Setup.
-Given a molecular graph $g^{(i)}\in\mathcal{G}=\{(g^{(i)},y^{(i)})\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as $f:g\rightarrow y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$.
+Given a dataset $\mathcal{G}=\{(g^{(i)},y^{(i)})\}$ consisting of molecular graphs $\{g^{(i)}\}$ with their property labels $\{y^{(i)}\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as denoted as $f:g\mapsto y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$.
 Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig.
@@
 3.1 Our Design: Lamole
-In this work, we use pre-trained Transformer-based LMs, e.g., Bert family models, as our backbone.
-Then, we fine-tune the LMs on the Group SELFIES corpus for explainable molecular property prediction.
+In this work, we pre-train Transformer-based LMs, e.g., Bert family models, using the Group SELFIES corpus to make the models understand the chemical semantics behind Group SELFIES strings.
+Then, we fine-tune the LMs with Group SELFIES strings and the molecular property labels for explainable molecular property prediction.
 An illustration of proposed Lamole is in Fig.
@@
 Information Flow-Based Explanations.
-To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature [13].
-However, we argue that using gradients alone cannot effectively capture functional group interactions.
+To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature (Selvaraju et al., 2017).
+However, we argue that using gradients alone cannot effectively capture substructure interactions.
 As depicted in Fig.
@@
 Firstly, we show the process of deriving the gradient-based explanations.
-Similar to GradCAM [13], the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by
-$\nabla\mathbf{h}_{j}^{(l)}=\frac{\partial\hat{y}}{\partial\mathbf{h}_{j}^{(l)}},$ (1)
-where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$.
+Similar to GradCAM (Selvaraju et al., 2017), the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by $\nabla\mathbf{h}_{j}^{(l)}=\partial\hat{y}/{\partial\mathbf{h}_{j}^{(l)}}$, where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$.
 Due to the skip connection in Fig.
-3 right panel, the input and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
+3 right panel, the input, and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
 $\mathbf{w}_{j}^{(l)}=\nabla\mathbf{h}_{j}^{(l)}\odot\mathbf{h}_{j}^{(l)},$ (1)
@@
 First, we formally define the plausibility of explanations.
-"Plausibility" refers to how the interpretation convinces humans [31, 32, 33].
+"Plausibility" refers to how the interpretation convinces humans (Wiegreffe & Pinter, 2019; Herman, 2017; Jacovi & Goldberg, 2020).
 Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
@@
 The higher the $\mathrm{EP}$ value, the greater the confidence of the explanation in matching the ground truth substructure.
+Eq.
+(4) defines the explanation plausibility based on the scores of two parts, i.e., the scores on ground truth and the scores on non-ground truth.
+The lower the scores on non-ground truth and the greater the scores on ground truth, the better explanation plausibility.
+Therefore, to maximize the plausibility, our objective can be transformed to minimize the importance score of non-ground truth and maximize the importance score of ground truth.
+To this end, we design a max-margin loss to optimize the importance score.
 In our work, the ground truth substructures are annotated by a binary mask vector $\mathbf{m}(g)\in\{0,1\}^{j}$.
@@
 The overall optimization objective of the fine-tuning stage is to minimize $\mathcal{L}_{CE}+\mathcal{L}_{M}$.
-By minimizing $\mathcal{L}_{M}$, Lamole aligns the explanations with the domain knowledge, enhancing the plausibility of explanations.
+The core of Eq.
+(5) is the discrepancy between the average importance score of ground truth and the average importance score of non-ground truth.
+By minimizing $\mathcal{L}_{M}$, the discrepancy between the two importance scores is maximized.
+In other words, the average importance score of non-ground truth is suppressed, and the average importance score of ground truth is increased.
+Finally, the explanation plausibility defined in Eq.
+(4) is improved.
 The next section will theoretically show that by using the designed marginal loss, the explanations can faithfully reflect the structure-property relationships.
@@
 Manifold Hypothesis.
-It is widely believed that natural data, including molecules, distribute around a manifold [34, 11, 35, 36].
-According to the manifold hypothesis for gradient-based explanations [34], if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained.
+It is widely believed that natural data, including molecules, distribute around a manifold (Bordt et al., 2023; Lin et al., 2022; Godwin et al., 2022; Singh et al., 2020).
+According to the manifold hypothesis for gradient-based explanations (Bordt et al., 2023), if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained.
 We call these features "causal features" in our work.
@@
 With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships.
-Due to the page limitation, we provide the proof in Appendix 1.
+Due to the page limitation, we provide the proof in Appendix.
 Theorem 1
@@
 We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction.
-Six mutagenicity datasets are Mutag [37], Mutagen [38], PTC-FM [39], PTC-FR [39], PTC-MM [39], and PTC-MR [39].
-For hepatotoxicity [39], the Liver dataset [40] is used.
+Six mutagenicity datasets are Mutag (Debnath et al., 1991), Mutagen (Morris et al., 2020), PTC-FM (Toivonen et al., 2003), PTC-FR (Toivonen et al., 2003), PTC-MM (Toivonen et al., 2003), and PTC-MR (Toivonen et al., 2003).
+For hepatotoxicity (Toivonen et al., 2003), the Liver dataset (Liu et al., 2015) is used.
 We used these datasets as the ground truth substructures in these datasets are known.
-For detailed information on these datasets, please refer to Appendix 2.
-Evaluation Metrics. We use three metrics for evaluating the performance of the proposed Lamole.
-  1. 1.
-    Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$.
-  2. 2.
-    Explanation Accuracy: We follow the experimental settings in GNNExplainer [12], which formulates the explanation problem as a binary classification of edges.
+For detailed information on these datasets and the ground truth substructures, please refer to Appendix.
+Evaluation Metrics.
+We use three metrics for evaluating the performance of the proposed Lamole.
+1) Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$.
+2) Explanation Accuracy: We follow the experimental settings in GNNExplainer (Ying et al., 2019), which formulates the explanation problem as a binary classification of edges.
 We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation.
 We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones.
-  3. 3.
-    Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
+3) Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
+Baselines.
+We combine Lamole into three BERT family models, such as DistilBert (Sanh, 2019), DeBerta (He et al., 2020), and Bert to evaluate the performance of the proposed Lamole.
+For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa (Chithrananda et al., 2020) and several GNNs including GCN (Kipf & Welling, 2016), DGCNN (Zhang et al., 2018), edGNN (Jaume et al., 2019), GIN (Xu et al., 2018), RW-GNN (Nikolentzos & Vazirgiannis, 2020), DropGNN (Papp et al., 2021), and IEGN (Maron et al., 2018).
+For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad (Smilkov et al., 2017), GradInput (Shrikumar et al., 2017), and GradCAM (Selvaraju et al., 2017), 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX (Lin et al., 2022), GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), and Gem (Lin et al., 2021).
+The details of experimental settings can be found in Appendix.
+4.2 Results
+Prediction Performance.
+Table 1 shows the classification accuracy of compared algorithms.
+As we can see, our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert not only can provide explainability but also can achieve comparable prediction accuracy as compared to existing predictive methods.
+In addition, Lamole models show superior performance over ChemBERTa.
+This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the chemical semantics and structure-property relationships.
 Table 1: Mean Classification Accuracy on the Seven Datasets (%)
 Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
-GCN [41] 84.6 78.9 54.8 63.0 57.8 53.3 41.1
-DGCNN [42] 85.8 74.8 57.3 63.5 61.0 58.6 44.6
-edGNN [43] 86.9 75.2 59.8 65.7 64.4 56.3 44.5
-GIN [44] 87.5 82.3 62.1 66.2 65.1 64.0 44.9
-RW-GNN [45] 87.2 80.3 61.9 64.0 62.4 57.0 43.2
-DropGNN [46] 89.4 80.7 62.0 66.0 63.7 64.2 45.0
-IEGN [47] 84.6 80.1 60.8 59.8 61.1 59.5 45.3
-ChemBERTa [10] 86.8 78.0 60.0 65.7 60.4 58.7 45.7
+GCN (Kipf & Welling, 2016) 84.6 78.9 54.8 63.0 57.8 53.3 41.1
+DGCNN (Zhang et al., 2018) 85.8 74.8 57.3 63.5 61.0 58.6 44.6
+edGNN (Jaume et al., 2019) 86.9 75.2 59.8 65.7 64.4 56.3 44.5
+GIN (Xu et al., 2018) 87.5 82.3 62.1 66.2 65.1 64.0 44.9
+RW-GNN (Nikolentzos & Vazirgiannis, 2020) 87.2 80.3 61.9 64.0 62.4 57.0 43.2
+DropGNN (Papp et al., 2021) 89.4 80.7 62.0 66.0 63.7 64.2 45.0
+IEGN (Maron et al., 2018) 84.6 80.1 60.8 59.8 61.1 59.5 45.3
+ChemBERTa (Chithrananda et al., 2020) 86.8 78.0 60.0 65.7 60.4 58.7 45.7
 Lamole+DistilBert 84.2 76.8 57.5 69.0 60.2 64.5 47.2
@@
 Lamole+Bert 88.2 74.5 62.4 70.0 61.2 66.0 47.5
-  • and denote the best result and second result, respectively.
-Baselines.
-We combine Lamole into three BERT family models, such as DistilBert [48], DeBerta [49], and Bert to evaluate the performance of the proposed Lamole.
-For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa [10] and several GNNs including GCN [41], DGCNN [42], edGNN [43], GIN [44], RW-GNN [45], DropGNN [46], and IEGN [47].
-For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad [50], CAM [51], GradInput [52], and GradCAM [13], 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX [11], GNNExplainer [12], PGExplainer [21], and Gem [20].
-Implementation Details.
-Following ChemBERTa [10], our used BERT family models are pre-trained on 100,000 molecules from the ZINC dataset [53].
-15% tokens in each input string are randomly for masked language learning.
-We finetuned the models on Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, PTC-MR, and Liver.
-For each dataset, the ratio of samples with annotation masks over the data size is 10%.
-The details of experimental settings can be found in Appendix 3.
-4.2 Results
+  • and denote the best and the second-best results, respectively.
+Explanation Performance.
+Table 2 presents the explanation accuracy of the compared explainability techniques.
+It should be noted that the ground truth annotations used in our work provide additional supervisory signals.
+Therefore, we also align the generated explanations of these baselines with the annotations for fair comparison, and the ground truth annotation rate is 10%.
+Lamole improves the explanation accuracy by $1.4\%\sim 14.3\%$ compared to the baseline methods.
+We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%).
+The impact of annotation rates can be found in Appendix.
+In addition, we discuss the rationale of using labeled annotations (see Appendix).
 Table 2: Mean Explanation Accuracy on the Seven Datasets (%)
 Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
-SmoothGrad+GCN [50] 68.0 66.1 66.8 62.2 64.5 61.9 66.0
-CAM+GCN [51] 66.8 65.4 66.9 62.7 61.4 65.5 67.5
-GradInput+GCN [52] 69.1 65.1 68.5 64.3 62.3 62.0 67.2
-GradCAM+GCN [13] 68.3 65.7 70.4 65.4 65.8 66.6 68.4
-SmoothGrad+Bert [50] 71.3 71.3 72.8 70.6 65.2 68.9 74.4
-CAM+Bert [51] 70.1 69.8 66.3 69.4 63.9 66.2 72.8
-GradInput+Bert [52] 73.1 71.1 72.1 66.6 64.0 68.1 75.0
-GradCAM+Bert [13] 73.5 71.5 76.3 69.0 69.1 71.2 75.8
-OrphicX [11] 76.4 70.4 71.1 68.4 69.0 69.1 73.0
-GNNExplainer [12] 69.5 63.0 68.6 67.3 66.0 66.3 71.4
-PGExplainer [21] 65.7 58.0 69.3 66.3 64.7 65.1 71.0
-Gem [20] 72.8 65.5 69.8 66.8 68.2 68.3 72,5
+GradInput+GCN (Shrikumar et al., 2017) 70.3 67.9 69.7 66.4 64.6 65.0 73.0
+GradCAM+GCN (Selvaraju et al., 2017) 69.8 67.0 71.0 67.9 66.2 67.3 69.4
+SmoothGrad+GCN (Smilkov et al., 2017) 69.2 66.8 67.5 62.6 64.9 63.1 66.4
+GradInput+Bert (Shrikumar et al., 2017) 75.1 72.6 73.0 68.9 65.6 69.6 75.6
+GradCAM+Bert (Selvaraju et al., 2017) 75.3 72.4 77.5 70.0 70.2 73.0 76.0
+SmoothGrad+Bert (Smilkov et al., 2017) 73.4 72.8 73.7 71.0 67.0 69.9 75.1
+GNNExplainer (Ying et al., 2019) 70.6 64.2 68.9 67.9 66.8 67.1 72.1
+PGExplainer (Luo et al., 2020) 66.5 58.7 70.3 68.0 65.9 67.0 71.5
+Gem (Lin et al., 2021) 73.7 66.0 71.3 69.0 68.9 69.2 73.6
+OrphicX (Lin et al., 2022) 78.0 71.5 74.6 70.4 70.9 71.4 74.0
 Lamole+DistilBert 70.9 73.0 74.0 70.2 69.6 78.1 76.1
@@
 Lamole+Bert 77.8 75.2 81.1 72.2 72.0 73.1 77.3
+  • and denote the best and the second-best results, respectively.
 Figure 4: Explanation visualization of one molecule (ID: 155) from the Mutag dataset.
@@
 (c) PTC-MR
-Figure 5: The explanation accuracy of Lamole with different annotation rates on the Mutag, PTC-FM, and PTC-MR datasets.
-(a) Mutag
-(b) PTC-FM
-(c) PTC-MR
 Figure 5: The explanation plausibility of the compared algorithms on the Mutag, PTC-FM, and PTC-MR datasets.
-Molecular Property Prediction.
-Table 1 shows the classification accuracy of compared algorithms.
-Our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert slightly outperform most compared algorithms on the Mutag, PTC-FM, PTC-FR, PTC-MR, and Liver datasets.
-Though the LMs, including Lamole and ChemBERTa, demonstrate good performance for molecular property prediction, they are inferior to GNN-based methods on the Mutagen and PTC-MM datasets.
-In addition, Lamole models show superior performance over ChemBERTa.
-This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the structure-property relationships.
-Explanation Performance.
-Table 2 presents the explanation accuracy of the compared explainability techniques.
-It is obvious that the LM-based methods using the Group SELFIES strings demonstrate significantly higher explanation accuracy over explainable GNNs.
-Specifically, Lamole improves the explanation accuracy by $1.4\%\sim 14.8\%$ compared to the baseline methods.
-This is because the molecular graph-based explainable GNNs can only highlight individual atoms and bonds rather than chemically meaningful substructures like functional groups.
-Instead, Group SELFIES explicitly encode the chemically meaningful substructures as tokens, and using Group SELFIES can help explainable methods to generate chemically meaningful explanations by highlighting the corresponding tokens.
-We also selected some representative molecules for explanation visualization.
+We selected some representative molecules for explanation visualization.
 These explanations are shown in Figs.
-1, 4, and 12$\sim$15 (see Appendix), respectively.
+1 and 4, respectively.
 The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments.
-As shown in Fig.
-4, our proposed Lamole assigns the highest two importance scores to the two ground truth function groups, i.e., benzene-1 and nitro group.
-In contrast, OrphicX cannot provide chemically meaningful explanations, and CradCAM+Bert cannot accurately identify the functional group interactions.
 Compared to baseline methods, Lamole provides chemically meaningful explanations.
 Particularly, the interaction among the functional groups is successfully captured.
-The visualization demonstrates Lamole’s superior interpretation in uncovering the structure-property relationships.
-We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%), as presented in Figs.
-5 and 10 (see Appendix).
-Clearly, using more molecules with ground truth annotations can continuously improve explanation accuracy.
-In addition, compared to the results without using the marginal loss $\mathcal{L}_{M}$, using only 10% molecules with ground truth annotations can significantly improve explanation accuracy by up to 5%, confirming the effectiveness of using annotations.
-Though increasing the annotated ground truth rate can only bring a limited increase in the explanation accuracy, there is a trade-off between the achieved explanation accuracy and the additional annotation costs.
+More visualization results can be found in Appendix.
 In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$.
 The statistical results of $\mathrm{EP}$ are presented in Figs.
-6 and 11 (see Appendix).
+5 and LABEL:epfr.
 From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property.
 Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$.
-The statistical results of $\mathrm{EP}$ in Figs.
-6 and 11 (see Appendix) indicate Lamole can provide explanations with higher confidence in matching the ground truth substructures.
-Due to the page limitation, more analysis regarding the ablation studies on attention weights and the marginal loss, the impact of annotation rates, the analysis regarding the explanation plausibility, and more visualization results can be found in the Appendix.
+More analysis regarding the explanation plausibility can be found in Appendix.
+4.3 Ablation Studies
+We conducted ablation studies for each component in our Lamole.
+Specifically, we removed the attention weights in the explanations, removed the marginal loss, and only used attention weights as explanations.
+The corresponding ablation algorithms are named Lamole ($-\alpha$), Lamole ($-\mathcal{L}_{M}$), and Lamole (Att), respectively.
+The results are shown in Fig.
+6.
+The explanation accuracy decreases by 1.4$\sim$2.3when removing the attention weights.
+Removing the marginal loss can decrease the explanation accuracy by 1.0$\sim$5.0%.
+Regarding the results of using only attention weights, The explanation accuracy decreases by 1.4$\sim$6.3%.
+The above results confirm the effectiveness of using the marginal loss, attention weights, and gradients.
+Figure 6: The explanation accuracy of Lamole+Bert, Lamole+Bert ($-\alpha$), Lamole+Bert ($-\mathcal{L}_{M}$), and Lamole (Att).
+From the perspective of model training, the marginal loss enables the model to be trained under the causal signals.
+To verify, we compared the classification performance with and without the marginal loss, as shown in Table 3.
+Without the marginal loss, the classification accuracy degrades by 0.7$\sim$3.7%.
+The above results indicate that marginal loss could help identify the causal features, thereby improving classification accuracy.
+Table 3: The classification performance with and without the marginal loss (%)
+Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
+Lamole-$\mathcal{L}_{M}$ 84.9 73.8 58.6 69.3 59.7 62.3 44.6
+Lamole 88.2 74.5 62.4 70.0 61.2 66.0 47.5
+To investigate the attention weights, the attention weights of two molecules are depicted in Fig.
+7.
+From Fig.
+7 middle panel, it can be found the correlation among the ground truth substructures is higher than others, showcasing the rationality of using attention weights to capture the functional group interactions.
+When we aggregate the attention weights for each token, as shown in Fig.
+7 right panel, the top two attention weights of the molecule (ID:155) match the two ground truth substructures.
+However, the top two attention weights of the molecule (ID:156) do not match the two ground truth substructures.
+The above results indicate that attention weights can capture the interactions and also confirm that "attention is not explanation" (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020).
+Due to the page limitation, the limitation of the proposed work is provided in Appendix.
+Figure 7: Attention weights visualization. The ground truth substructures are highlighted in red.
 5 Conclusions
-This work developed a new framework for explainable molecular property prediction based on language models, called Lamole.
-Lamole uses Group SELFIES as input for perceiving the chemically meaningful semantics.
-By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations that can capture the functional group interactions.
-Furthermore, one marginal loss is designed to calibrate the explanations by aligning them with the chemists’ annotation.
-With the above designs, Lamole can provide chemically meaningful explanations while faithfully uncovering the structure-property relationships.
+This work proposed Lamole for explainable molecular property prediction based on language models.
+Lamole uses Group SELFIES as input for chemically meaningful semantics.
+By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations to quantify each chemically meaningful substructure’s impact on the model’s output.
+Furthermore, one marginal loss is designed to calibrate the explanations to be more faithful by aligning them with the chemists’ annotation.
 Lamole ’s effectiveness has been demonstrated through theoretical analysis and extensive experimental validation.

### v4

@@
 Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships.
-In this work, we develop a framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole..
+In this work, we develop a framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations.
 We take a string-based molecular representation — Group SELFIES — as input tokens to pre-train and fine-tune our Lamole, as it provides chemically meaningful semantics.
 By disentangling the information flows of Lamole, we propose considering both self-attention weights and gradients for better quantification of each chemically meaningful substructure’s impact on the model’s output.
-To make the explanations more faithfully respect the structure-property relationship, we then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations.
+To make the explanations more faithful to the structure-property relationship, we then carefully craft a marginal loss to explicitly optimize the explanations to align with the chemists’ annotations.
 We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations.
-Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.3\%$, being the state-of-the-art in explainable molecular property prediction.
-1 Introduction
-Molecular property prediction aims to reveal the molecular structures-property relationships, assisting scientists in screening molecules for various applications such as drug discovery and material design (Fang et al., 2022; Deng et al., 2023; Tripp et al., 2023; Wang et al., 2024; Ekström Kelvinius et al., 2023; Hong et al., 2024).
+Experimental results over eight datasets demonstrate Lamole can achieve comparable prediction accuracy and boost the explanation accuracy by up to $14.3\%$, being the state-of-the-art in explainable molecular property prediction.
+To further illustrate the actionable utility of the explanations derived from Lamole, we integrated the framework with an evolutionary algorithm.
+This integration established an interpretable optimization pipeline for molecular editing, demonstrating that Lamole functions beyond simple post-hoc analysis but serves as a practical guide for molecule discovery.
+I Introduction
+Molecular property prediction aims to reveal the structures-property relationships, assisting scientists in screening molecules for various applications such as drug discovery and material design [53, 12, 10, 51, 11, 17, 24, 26, 27, 25].
 Several learning-based models are devised based on the underlying molecular representations, such as graph-based and string-based molecular representations.
-Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES (Weininger, 1988)), stand out for their simplicity and adaptability (Deng et al., 2023; Wigh et al., 2022; Cheng et al., 2023).
-By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert (Kenton & Toutanova, 2019) offer higher throughput and accuracy for molecular property prediction (Deng et al., 2023; Chithrananda et al., 2020).
+Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES [55]), stand out for their simplicity and adaptability [10, 58, 6].
+By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert [22] offer higher throughput and accuracy for molecular property prediction [10, 7].
 Figure 1: (a) The molecule visualization of prediction/explanation.
 The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule.
-(b)-(e) are the explanation results obtained with various methods: (b) OrphicX (Lin et al., 2022); (c) GNNExplainer (Ying et al., 2019), (d) GNN with gradient-based explainability technique (GradCAM (Selvaraju et al., 2017)); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES (Cheng et al., 2023) as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
-Figure 2: The SMILES and Group SELFIES strings of p-nitrobenzoic acid molecule ($\mathrm{C_{7}H_{5}NO_{4}}$): The tokens in the Group SELFIES string highlighted by color are the corresponding functional groups.
+(b)-(e) are the explanation results obtained with various methods: (b) OrphicX [30]; (c) GNNExplainer [64], (d) GNN with gradient-based explainability technique (GradCAM [44]); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES [6] as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
 Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains.
 Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions.
-These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design (Wu et al., 2023; Wellawatte et al., 2023; Das et al., 2022).
-With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction (Proietti et al., 2024; Ying et al., 2019; Ye et al., 2023; Lin et al., 2021).
-However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously.
-Specifically, 1) from the molecular representation perspective, the commonly used representations do not explicitly encode the chemically meaningful substructures; current explainability methods can only highlight individual atoms and bonds as explanations (see Fig.
-1 (b)$\sim$(e)).
-2) From the perspective of explainability techniques, current methods suffer from two main limitations.
-First, they cannot effectively capture the interactions between functional groups within the molecular structure.
-Second, they could not generate explanations that align with chemists’ intuition.
-As a result, they fail to produce explanations that faithfully reflect the structure-property relationships.
-(see Fig.
-1 (f)).
-Therefore, an effective framework is imperative for explainable molecule property predictions.
-The recently proposed string-based molecular representation — Group SELFIES (Cheng et al., 2023) — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations.
+These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design [60, 56, 8].
+With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction [39, 64, 63, 29].
+[…]
+The recently proposed string-based molecular representation — Group SELFIES [6] — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations.
 As shown in Fig.
@@
   1. 1.
-    We found that existing explainability techniques fail to provide chemically meaningful explanations, perceive functional group interactions, and reveal molecular structure-property relationships faithfully.
-To address the issues, we use Group SELFIES to pre-train and fine-tune LMs to make LMs easily understand the chemically meaningful semantics.
-Moreover, the process of generating explanations should reflect the reasoning process behind the model architecture.
-Therefore, by disentangling the information flows of Transformer-based LMs, we integrate the self-attention weights and gradients to capture the substructure interactions to better quantify each chemically meaningful substructure’s contribution to the predicted molecular properties.
+    We found that existing explainability techniques fail to reveal molecular structure-property relationships faithfully.
+Therefore, to provide explanations directly aligned with chemical intuition, we first leverage the Group SELFIES representation—which inherently encodes chemically meaningful substructures—for both pre-training and fine-tuning our LMs.
+By integrating this chemically grounded semantic input, our framework enables LMs to directly comprehend and operate on functional group semantics, moving beyond low-level features to generate explanations that are immediately relevant and intuitive for chemists.
   2. 2.
-    To make the explanations more faithfully respect the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations.
-We show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%.
+    We analyzed the limitations of existing methods in revealing molecular structure-property relationships, noting that approaches relying solely on attention or gradients as explanations often fail to adequately reflect the model’s underlying reasoning process.
+For an explanation to be faithful, the generation process must be closely coupled with the mechanism of the model architecture.
+Therefore, by carefully disentangling the information flows within Transformer-based LMs, we propose an information flow-based explainable method.
+This strategy is designed specifically to capture the interactions between chemically meaningful substructures, thereby providing a more accurate quantification of each substructure’s contribution to the predicted molecular properties.
   3. 3.
-    We first bridge the manifold hypothesis with explainable molecular property prediction.
+    To make the explanations more faithful to the structure-property relationships, we elaborate on one marginal loss to calibrate the explanations by aligning them with the chemists’ annotations.
+Empirically, we show that using only a few molecules with ground truth annotations can significantly improve the explanation accuracy by up to 5%.
+Furthermore, theoretically, we first bridge the manifold hypothesis with explainable molecular property prediction.
 We theoretically demonstrate that the elaborated marginal loss aligns explanations with the data manifold, respecting the structure-property relationship.
-Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.3\%$.
-We also quantitatively evaluate explanations based on the first proposed plausibility metric.
-Compared to alternative baselines, the explanation plausibility of Lamole is improved by up to $9\%$.
-Extensive experimental studies demonstrate Lamole achieves state-of-the-art performance in explainable molecular property prediction.
-2 Related Work
-Several explainable GNNs are proposed to explain the relationship between the input graph and the prediction (Sun et al., 2023; Xiong et al., 2019; Lin et al., 2022; Ying et al., 2019; Luo et al., 2020; Lin et al., 2021).
-Among these works, structure similarity or attention weights are proposed to capture structural interaction (Sun et al., 2023; Xiong et al., 2019).
-However, two similar substructures do not necessarily lead to interaction between them, and attention weights are often inconsistent with the feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020).
-In addition, as shown in Fig.
-1 (b)$\sim$(d), some trivial structures received relatively high importance scores, indicating the explanations might not align well with the chemical concepts.
-On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction (Chithrananda et al., 2020; Wang et al., 2019; Ahmad et al., 2022; Ross et al., 2022).
+  4. 4.
+    We further demonstrate the actionable utility of the explanations derived from Lamole through a dedicated molecular editing case study.
+By integrating Lamole’s identified key functional groups with an evolutionary algorithm, we establish an interpretable optimization pipeline that guides the molecular optimization process to enhance target properties.
+This integration showcases Lamole’s capability to bridge predictive insights with rational molecular design and optimization, paving the way for more efficient chemical discovery.
+  5. 5.
+    We evaluated our method across eight diverse datasets spanning both classification and regression tasks, including a wide range of molecular properties such as hepatotoxicity, mutagenicity, water solubility, and permeability.
+Experimental results demonstrate that our proposed Lamole achieves comparable predictive accuracy relative to other established baselines while exhibiting superior explanation accuracy, confirming the broad applicability of Lamole across different property prediction tasks.
+Figure 2: The SMILES and Group SELFIES strings of p-nitrobenzoic acid molecule ($\mathrm{C_{7}H_{5}NO_{4}}$): The tokens in the Group SELFIES string highlighted by color are the corresponding functional groups.
+II Related Work
+Molecular property prediction is a fundamental task aimed at elucidating the relationship between molecular structure and function.
+Various learning-based models have been developed based on distinct molecular representations, primarily categorized into graph-based and string-based formats.
+GNNs explicitly model molecular systems by incorporating atomic positions and pairwise interactions through message passing.
+For instance, SchNOrb [43] adapted distance-based convolution by integrating symmetry-adapted atomic orbital features for electronic structure prediction.
+MiCaM [13] introduced a fragment-based deep generative model that enhances novelty by automatically discovering connection-aware motifs through iterative subgraph merging, which are then utilized in a dedicated generator for de novo molecular generation.
+Alternatively, string-based molecular representations, such as SMILES, are favored for their simplicity and adaptability within sequential models.
+TamGen [59], for example, leveraged a GPT-like chemical language model to achieve target-aware molecule generation and compound refinement.
+This work demonstrated the efficacy of GPT-like architectures and training strategies for generation within the one-dimensional SMILES chemical space by integrating specific modules for encoding protein structure and seeding compound information.
+[…]
+On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction [7, 52, 4, 40].
 However, the "black-box" characteristics of LMs hamper trust use of these potent computational tools in scientific domains.
@@
 One way is to use the attention weights over the input tokens.
-However, recent studies suggest that "attention is not explanation" because attention weights could not reflect the true feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020).
+However, recent studies suggest that "attention is not explanation" because attention weights could not reflect the true feature importance [20, 45, 1].
 Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the input importance.
-However, the generated explanations may change drastically with very small perturbations (Agarwal et al., 2021).
-Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature (Selvaraju et al., 2017).
-However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations (Adebayo et al., 2018; Agarwal et al., 2021; Rudin, 2019).
+However, the generated explanations may change drastically with very small perturbations [3].
+Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature [44].
+However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations [2, 3, 41].
 Therefore, the explanation generation process should reflect the model reasoning process behind the model architectures.
 To this end, we disentangle the model architectures’ information flows to generate explanations that faithfully reveal the structure-property relationship.
-3 Methodology
+III Methodology
 Problem Setup.
 Given a dataset $\mathcal{G}=\{(g^{(i)},y^{(i)})\}$ consisting of molecular graphs $\{g^{(i)}\}$ with their property labels $\{y^{(i)}\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as denoted as $f:g\mapsto y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$.
-Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig.
+Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., $\cdot$] in Fig.
 2.
@@
 In addition, a marginal loss $\mathcal{L}_{M}$ is designed to align explanations with the chemists’ annotations $\mathbf{m}$.
-3.1 Our Design: Lamole
+III-A Our Design: Lamole
 In this work, we pre-train Transformer-based LMs, e.g., Bert family models, using the Group SELFIES corpus to make the models understand the chemical semantics behind Group SELFIES strings.
@@
 In what follows, we will introduce the detailed design of Lamole.
-Fine-Tuning Stage.
-Fig.
-3 shows the fine-tuning stage of the proposed Lamole.
-We assume that the Transformer encoder in Lamole stacks $L$ identical Transformer layers to encode the molecular string $s$ as token embedding $\mathbf{h}^{(l)}=\{\mathbf{h}_{1}^{(l)},\ldots,\mathbf{h}_{j}^{(l)}\}$, where $\mathbf{h}^{(l)}$ is the token embedding at the $l$-th layer.
-We use the self-attention weighted average embedding $\mathbf{h}_{o}=(\sum_{j=1}\alpha_{j}\cdot\mathbf{h}_{j}^{(L)})/(\sum_{j=1}\alpha_{j})$ for molecular property prediction, where $\alpha_{j}$ is the attention weight of the $j$-th token.
-A multilayer perceptron (MLP) classifier is added to predict the molecular property $\hat{y}=\mathrm{MLP}(\mathbf{h}_{o})$ by minimizing the classification cross-entropy loss $\mathcal{L}_{CE}(y,\hat{y})$.
+[…]
+Analysis of Foundational Explanation Methods.
+Before introducing our information flow-based approach, we first analyze two foundational methods—attention and gradients—and detail their limitations when applied to Transformer-based LMs.
+1) Attention as Explanations: This approach quantifies the importance of each molecule token to the LM’s final decision.
+This is achieved by extracting the softmax cross-attention from the starting answer token to all input tokens across all layers of the backbone LM.
+The resulting prediction-to-token attention is represented as $A_{\text{st}}(s)\in\mathbb{R}^{L\times H\times 1\times T}$, where $L$ and $H$ denote the number of layers and heads-per-layer, respectively, and $T$ is the total number of tokens provided.
+We subsequently compute the prediction-to-token attention $\hat{A}_{\text{st}}(s)$ by averaging across attention heads: $\hat{A}_{\text{st}}(s)=\frac{1}{H}\sum_{h=1}^{H}A_{\text{st}}(s)$.
+However, recent studies suggest that “attention is not explanation" because the attention scores often fail to reflect the true importance of the input features to the model’s decision [20, 45, 1].
+2).
+Gradients as Explanations: This common practice utilizes gradient-based methods to determine feature importance by analyzing the partial derivatives of the output with respect to each input feature [44].
+We seek to identify relevant tokens by computing the gradient of the LM’s decision with respect to the input tokens.
+The pure gradient is versatile as it is not restricted to a Transformer architecture.
+Specifically, for a molecule-label pair $(s,y)$, the importance is computed as $G(s,y)=\|\nabla_{s}v(s,y)\|_{2}$, where $v(s,y)$ is the logarithm of the maximum output probability of the LM.
+While both practices are widely used to derive explanations, we argue that relying on gradients alone is insufficient for effectively capturing substructure interactions.
+As illustrated in Fig.
+1 (f), a pure gradient-based method like GradCAM can inaccurately attribute a property to an unrelated structural fragment, such as incorrectly highlighting a distant benzene ring alongside the true nitro group contributor.
+The primary reason for this deficiency lies in the internal architecture of the Transformer.
+As shown in the right panel of Fig.
+3, the output is influenced by a confluence of factors, including attention weights, gradients, and the original input itself, primarily due to the ubiquitous skip connections and the attention mechanism.
+Consequently, using gradients in isolation fails to capture the rich network of information exchange—the interactions—that occurs between substructures via the attention layers.
+To address this limitation, we leverage both attention and gradients, integrating them alongside the input to derive a more comprehensive explanation.
+The following section details this integration of attention weights into our gradient-based explanation framework.
 Information Flow-Based Explanations.
-To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature (Selvaraju et al., 2017).
-However, we argue that using gradients alone cannot effectively capture substructure interactions.
-As depicted in Fig.
-1 (f), the gradient-based method CradCAM incorrectly attributes the property to the nitro group and a benzene ring that is not connected to the nitro group.
-We illustrate the possible reason by disentangling the information flows of the Transformer.
-As shown in Fig.
-3 right panel, due to the skip connection and attention mechanism, both attention weights, gradients, and the input contribute to the outputs.
-Therefore, using gradients alone as explanations could fall short of capturing interactions.
-To address the issue, we leverage both attention and gradients, as well as the input, to derive explanations.
-Below, we will elaborate on integrating attention weights into gradient-based explanations.
 Firstly, we show the process of deriving the gradient-based explanations.
-Similar to GradCAM (Selvaraju et al., 2017), the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by $\nabla\mathbf{h}_{j}^{(l)}=\partial\hat{y}/{\partial\mathbf{h}_{j}^{(l)}}$, where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$.
+Similar to GradCAM [44], the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by $\nabla\mathbf{h}_{j}^{(l)}=\partial\hat{y}/{\partial\mathbf{h}_{j}^{(l)}}$, where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$.
 Due to the skip connection in Fig.
@@
 First, we formally define the plausibility of explanations.
-"Plausibility" refers to how the interpretation convinces humans (Wiegreffe & Pinter, 2019; Herman, 2017; Jacovi & Goldberg, 2020).
+"Plausibility" refers to how the interpretation convinces humans [57, 16, 19].
 Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
+Eq.
+(4) defines the explanation plausibility based on the scores of two parts, i.e., the scores on ground truth and the scores on non-ground truth.
+The lower the scores on non-ground truth and the greater the scores on ground truth, the better the explanation plausibility.
+Therefore, to maximize the plausibility, our objective can be transformed to minimize the importance score of non-ground truth and maximize the importance score of ground truth.
 Definition 1
@@
 The higher the $\mathrm{EP}$ value, the greater the confidence of the explanation in matching the ground truth substructure.
-Eq.
-(4) defines the explanation plausibility based on the scores of two parts, i.e., the scores on ground truth and the scores on non-ground truth.
-The lower the scores on non-ground truth and the greater the scores on ground truth, the better explanation plausibility.
-Therefore, to maximize the plausibility, our objective can be transformed to minimize the importance score of non-ground truth and maximize the importance score of ground truth.
 To this end, we design a max-margin loss to optimize the importance score.
@@
 To achieve the goal, a max-margin loss is designed by maximizing the mean value of the importance scores of tokens that have mask values of 1 while minimizing the mean value of importance scores for tokens with mask values of 0.
-$\mathcal{L}_{M}(\mathbf{v},\mathbf{m})=\mathbb{E}_{g\in\mathcal{G}}\left[\max\left(0,\frac{\sum_{j=1}(1-\mathbf{m}_{j}(g))\cdot\mathbf{v}_{j}(g)}{N_{s}}-\frac{\sum_{j=1}\mathbf{m}_{j}(g)\cdot\mathbf{v}_{j}(g)}{N_{c}}\right)+\triangle_{1}\right],$ (5)
+$\mathcal{L}_{M}(\mathbf{v},\mathbf{m})=\mathbb{E}_{g\in\mathcal{G}}\Bigg[\max\Bigg(0,\frac{\sum_{j=1}(1-\mathbf{m}_{j}(g))\cdot\mathbf{v}_{j}(g)}{N_{s}}\\ -\frac{\sum_{j=1}\mathbf{m}_{j}(g)\cdot\mathbf{v}_{j}(g)}{N_{c}}\Bigg)+\triangle_{1}\Bigg],$ (5)
 where $\triangle_{1}$ is a margin term, $N_{s}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 0, and $N_{c}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 1.
@@
 The next section will theoretically show that by using the designed marginal loss, the explanations can faithfully reflect the structure-property relationships.
-3.2 Theoretical Analysis
+III-B Theoretical Analysis
 We bridge the manifold hypothesis with the marginal loss to theoretically show that the explanations can respect the structure-property relationships.
@@
 Manifold Hypothesis.
-It is widely believed that natural data, including molecules, distribute around a manifold (Bordt et al., 2023; Lin et al., 2022; Godwin et al., 2022; Singh et al., 2020).
-According to the manifold hypothesis for gradient-based explanations (Bordt et al., 2023), if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained.
+It is widely believed that natural data, including molecules, distribute around a manifold [5, 30, 14, 47].
+According to the manifold hypothesis for gradient-based explanations [5], if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained.
 We call these features "causal features" in our work.
@@
 With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships.
-Due to the page limitation, we provide the proof in Appendix.
 Theorem 1
@@
 (5) aligns the gradient-based explanations $\nabla_{g}\log p(y|g)$ with the tangent space of the causal feature manifold $\mathcal{M}_{c}$, thus respecting the structure-property relationships.
-4 Experiments
-4.1 Experimental Setup
+Proof of Theorem 1: The gradient with respect to the prediction $\nabla_{g}\log p(y|g)$ can be decomposed into the gradient on the causal features and spurious features, respectively,
+$\nabla_{g}\log p(y|g)=\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)\\ +\nabla_{g}\log p(y|g)\odot(1-\mathbf{m}(g)).$ (7)
+By minimizing the loss of Eq.
+(5), the gradients on spurious features $\nabla_{g}\log p(y|g)\odot(1-\mathbf{m}(g))$ are suppressed, and $\nabla_{g}\log p(y|g)$ approximates $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$.
+Therefore, we have $\nabla_{g}\log p(y|g)\approx\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$.
+On the other hand, $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$ can be rewritten as
+$\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)=\nabla_{g}\log p(g|y)\odot\mathbf{m}(g)\\ -\sum_{j}p(y=j|g)\nabla_{g}\log p(g|y=j)\odot\mathbf{m}(g).$ (8)
+Because the data distribution $p(g|y)\odot\mathbf{m}(g)$ reflects the causal feature manifold $\mathcal{M}_{c}$, the gradient of the distribution $\nabla_{g}p(g|y)\odot\mathbf{m}(g)$ represents the tangent space of the causal feature manifold $\mathcal{M}_{c}$.
+In addition, Eq.
+(8) shows that the $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$ is a linear combination of $\nabla_{g}p(g|y)\odot\mathbf{m}(g)$, so $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$ also lies tangent space of the manifold $\mathcal{M}_{c}$.
+Together with Eq.
+(7) and Eq.
+(8), we prove the gradient-based explanations $\nabla_{g}\log p(y|g)$ lies tangent space of the manifold $\mathcal{M}_{c}$.
+This indicates by minimizing the loss of Eq.
+(5), the model $p(y|g)$ has reflected the causal feature manifold.
+According to the manifold hypothesis, the features on the causal feature manifold contribute to the molecular property.
+Therefore, the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the causal features, thus revealing the structure-property relationships.
+This completes the proof.
+IV Experiments
+IV-A Experimental Setup
 Datasets.
 We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction.
-Six mutagenicity datasets are Mutag (Debnath et al., 1991), Mutagen (Morris et al., 2020), PTC-FM (Toivonen et al., 2003), PTC-FR (Toivonen et al., 2003), PTC-MM (Toivonen et al., 2003), and PTC-MR (Toivonen et al., 2003).
-For hepatotoxicity (Toivonen et al., 2003), the Liver dataset (Liu et al., 2015) is used.
-We used these datasets as the ground truth substructures in these datasets are known.
-For detailed information on these datasets and the ground truth substructures, please refer to Appendix.
+Six mutagenicity datasets are Mutag [9], Mutagen [35], PTC-FM [50], PTC-FR [50], PTC-MM [50], and PTC-MR [50].
+For hepatotoxicity [50], the Liver dataset [31] is used.
+Larger-sized molecules typically include more complex structures.
+The datasets that we used contained relatively large molecules.
+The maximal number of atoms of Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, PTC-MR, and Liver are 26, 417, 64, 64, 64, 64, and 157, respectively.
+The details of the used dataset are provided in Table I.
+TABLE I: Statistiscal Information of the Datasets
+Datasets Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
+Graphs 188 4337 349 351 336 344 587
+Classes 2 2 2 2 2 2 3
+Max nodes 26 417 64 64 64 64 157
+Avg nodes 17.9 29 14.1 14.6 14 14.3 25.6
+Avg edges 19.8 30 14.5 15 14.3 14.7 27.4
+Ground truth 120 724 58 49 51 61 187
+  • denotes the number of molecules with known ground truth substructures.
+Following OrphicX [30], on the Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, and PTC-MR datasets, we only consider the explanations for the mutagenic class, because the molecules of the non-mutagenic class have no ground truth.
+Although some works used single $\mathrm{N}=\mathrm{N}$, $\mathrm{NO_{2}}$, or $\mathrm{NH_{2}}$ as ground truth, this is not reasonable, as 32% of non-mutagenic graphs in Mutagen containing at least single $\mathrm{NO_{2}}$ or $\mathrm{NH_{2}}$.
+In fact, the ground truth for the mutagenic class is the benzene with a chemical group on it, such as $\mathrm{N}=\mathrm{N}$, $\mathrm{NO_{2}}$, and $\mathrm{NH_{2}}$ [29, 30, 38].
+For the Liver dataset, the molecules of possible hepatotoxicity with ground truth substructures and hepatotoxicity with ground truth substructures are collected for explainable molecular property prediction.
+The twelve ground truth substructures of the Liver dataset are shown in Fig.
+4.
+To assess the generalizability of our method’s explainable prediction ability, we performed an extended evaluation on two critical absorption, distribution, metabolism, and excretion (ADME) properties: solubility and permeability.
+The evaluation was conducted using the ZINC dataset [18].
+Lacking expert-annotated references for these complex endpoints, we employed an automated, descriptor-driven approach to generate the required ground-truth functional group labels.
+The annotation procedure for the ground truth substructures in the dataset is detailed in the supplemental material.
+Figure 4: Twelve ground truth substructures of the Liver dataset.
+Lowercase element symbols represent aromatic atoms of the element; the letter "a" matches any aromatic atom.
+Elements in square brackets match any of the elements in a molecule.
 Evaluation Metrics.
-We use three metrics for evaluating the performance of the proposed Lamole.
+We use four metrics for evaluating the performance of the proposed Lamole.
 1) Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$.
-2) Explanation Accuracy: We follow the experimental settings in GNNExplainer (Ying et al., 2019), which formulates the explanation problem as a binary classification of edges.
+2) Explanation Accuracy: We follow the experimental settings in GNNExplainer [64], which formulates the explanation problem as a binary classification of edges.
 We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation.
 We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones.
-3) Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
+3) Explanations’ Fidelity: If an explanation is faithful, removing the most important fragments identified can cause a significant change in the model’s prediction score for the target property.
+The definition is as follows: $\text{Fidelity}(G,G_{\text{ex}})=[f(G)]_{c^{*}}-[f(G/G_{\text{ex}})]_{c^{*}}$, where $G$ is the original molecular inputs, $G_{\text{ex}}$ is the set of key functional groups/fragments identified, and $f(\cdot)$ is the model’s classification probability or regression output.
+4) Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
 Baselines.
-We combine Lamole into three BERT family models, such as DistilBert (Sanh, 2019), DeBerta (He et al., 2020), and Bert to evaluate the performance of the proposed Lamole.
-For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa (Chithrananda et al., 2020) and several GNNs including GCN (Kipf & Welling, 2016), DGCNN (Zhang et al., 2018), edGNN (Jaume et al., 2019), GIN (Xu et al., 2018), RW-GNN (Nikolentzos & Vazirgiannis, 2020), DropGNN (Papp et al., 2021), and IEGN (Maron et al., 2018).
-For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad (Smilkov et al., 2017), GradInput (Shrikumar et al., 2017), and GradCAM (Selvaraju et al., 2017), 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX (Lin et al., 2022), GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), and Gem (Lin et al., 2021).
-The details of experimental settings can be found in Appendix.
-4.2 Results
-Prediction Performance.
-Table 1 shows the classification accuracy of compared algorithms.
-As we can see, our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert not only can provide explainability but also can achieve comparable prediction accuracy as compared to existing predictive methods.
-In addition, Lamole models show superior performance over ChemBERTa.
-This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the chemical semantics and structure-property relationships.
-Table 1: Mean Classification Accuracy on the Seven Datasets (%)
-Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
-GCN (Kipf & Welling, 2016) 84.6 78.9 54.8 63.0 57.8 53.3 41.1
-DGCNN (Zhang et al., 2018) 85.8 74.8 57.3 63.5 61.0 58.6 44.6
-edGNN (Jaume et al., 2019) 86.9 75.2 59.8 65.7 64.4 56.3 44.5
-GIN (Xu et al., 2018) 87.5 82.3 62.1 66.2 65.1 64.0 44.9
-RW-GNN (Nikolentzos & Vazirgiannis, 2020) 87.2 80.3 61.9 64.0 62.4 57.0 43.2
-DropGNN (Papp et al., 2021) 89.4 80.7 62.0 66.0 63.7 64.2 45.0
-IEGN (Maron et al., 2018) 84.6 80.1 60.8 59.8 61.1 59.5 45.3
-ChemBERTa (Chithrananda et al., 2020) 86.8 78.0 60.0 65.7 60.4 58.7 45.7
+We combine Lamole into three BERT family models, such as DistilBert [42], DeBerta [15], and Bert to evaluate the performance of the proposed Lamole.
+For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa [7] and several GNNs including GCN [23], DGCNN [66], edGNN [21], GIN [62], RW-GNN [36], DropGNN [37], and IEGN [34].
+For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad [48], GradInput [46], and GradCAM [44], 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX [30], GNNExplainer [64], PGExplainer [33], and Gem [29].
+The details of parameter settings can be found in the supplemental material.
+TABLE II: Mean Classification Accuracy on the Seven Datasets (%)
+Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
+GCN [23] 84.6 78.9 54.8 63.0 57.8 53.3 41.1
+DGCNN [66] 85.8 74.8 57.3 63.5 61.0 58.6 44.6
+edGNN [21] 86.9 75.2 59.8 65.7 64.4 56.3 44.5
+GIN [62] 87.5 82.3 62.1 66.2 65.1 64.0 44.9
+RW-GNN [36] 87.2 80.3 61.9 64.0 62.4 57.0 43.2
+DropGNN [37] 89.4 80.7 62.0 66.0 63.7 64.2 45.0
+IEGN [34] 84.6 80.1 60.8 59.8 61.1 59.5 45.3
+ChemBERTa [7] 86.8 78.0 60.0 65.7 60.4 58.7 45.7
 Lamole+DistilBert 84.2 76.8 57.5 69.0 60.2 64.5 47.2
@@
   • and denote the best and the second-best results, respectively.
-Explanation Performance.
-Table 2 presents the explanation accuracy of the compared explainability techniques.
-It should be noted that the ground truth annotations used in our work provide additional supervisory signals.
-Therefore, we also align the generated explanations of these baselines with the annotations for fair comparison, and the ground truth annotation rate is 10%.
-Lamole improves the explanation accuracy by $1.4\%\sim 14.3\%$ compared to the baseline methods.
-We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%).
-The impact of annotation rates can be found in Appendix.
-In addition, we discuss the rationale of using labeled annotations (see Appendix).
-Table 2: Mean Explanation Accuracy on the Seven Datasets (%)
-Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
-GradInput+GCN (Shrikumar et al., 2017) 70.3 67.9 69.7 66.4 64.6 65.0 73.0
-GradCAM+GCN (Selvaraju et al., 2017) 69.8 67.0 71.0 67.9 66.2 67.3 69.4
-SmoothGrad+GCN (Smilkov et al., 2017) 69.2 66.8 67.5 62.6 64.9 63.1 66.4
-GradInput+Bert (Shrikumar et al., 2017) 75.1 72.6 73.0 68.9 65.6 69.6 75.6
-GradCAM+Bert (Selvaraju et al., 2017) 75.3 72.4 77.5 70.0 70.2 73.0 76.0
-SmoothGrad+Bert (Smilkov et al., 2017) 73.4 72.8 73.7 71.0 67.0 69.9 75.1
-GNNExplainer (Ying et al., 2019) 70.6 64.2 68.9 67.9 66.8 67.1 72.1
-PGExplainer (Luo et al., 2020) 66.5 58.7 70.3 68.0 65.9 67.0 71.5
-Gem (Lin et al., 2021) 73.7 66.0 71.3 69.0 68.9 69.2 73.6
-OrphicX (Lin et al., 2022) 78.0 71.5 74.6 70.4 70.9 71.4 74.0
+TABLE III: Mean Explanation Accuracy on the Seven Datasets (%)
+Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
+GradInput+GCN [46] 70.3 67.9 69.7 66.4 64.6 65.0 73.0
+GradCAM+GCN [44] 69.8 67.0 71.0 67.9 66.2 67.3 69.4
+SmoothGrad+GCN [48] 69.2 66.8 67.5 62.6 64.9 63.1 66.4
+GradInput+Bert [46] 75.1 72.6 73.0 68.9 65.6 69.6 75.6
+GradCAM+Bert [44] 75.3 72.4 77.5 70.0 70.2 73.0 76.0
+SmoothGrad+Bert [48] 73.4 72.8 73.7 71.0 67.0 69.9 75.1
+GNNExplainer [64] 70.6 64.2 68.9 67.9 66.8 67.1 72.1
+PGExplainer [33] 66.5 58.7 70.3 68.0 65.9 67.0 71.5
+Gem [29] 73.7 66.0 71.3 69.0 68.9 69.2 73.6
+OrphicX [30] 78.0 71.5 74.6 70.4 70.9 71.4 74.0
 Lamole+DistilBert 70.9 73.0 74.0 70.2 69.6 78.1 76.1
@@
   • and denote the best and the second-best results, respectively.
-Figure 4: Explanation visualization of one molecule (ID: 155) from the Mutag dataset.
+TABLE IV: MEAN EXPLANATIONS’ FIDELITY ON THE SEVEN DATASETS
+Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
+SmoothGrad+GCN [48] 0.31 0.28 0.30 0.25 0.29 0.27 0.33
+CAM+GCN [67] 0.30 0.27 0.29 0.26 0.28 0.32 0.35
+GradInput+GCN [46] 0.32 0.26 0.31 0.28 0.25 0.26 0.34
+GradCAM+GCN [44] 0.33 0.29 0.34 0.30 0.31 0.33 0.36
+SmoothGrad+Bert [48] 0.40 0.38 0.42 0.39 0.35 0.38 0.45
+CAM+Bert [67] 0.39 0.37 0.36 0.38 0.34 0.37 0.43
+GradInput+Bert [46] 0.42 0.40 0.41 0.36 0.35 0.39 0.46
+GradCAM+Bert [44] 0.43 0.41 0.47 0.40 0.42 0.44 0.48
+OrphicX [30] 0.45 0.43 0.44 0.40 0.41 0.43 0.47
+GNNExplainer [64] 0.35 0.31 0.34 0.32 0.33 0.35 0.40
+PGExplainer [33] 0.33 0.29 0.35 0.31 0.32 0.34 0.39
+Gem [29] 0.41 0.36 0.38 0.37 0.39 0.40 0.44
+Lamole+DistilBert 0.50 0.52 0.55 0.51 0.53 0.62 0.58
+Lamole+DeBerta 0.58 0.60 0.65 0.55 0.56 0.61 0.59
+Lamole+Bert 0.61 0.62 0.68 0.57 0.58 0.57 0.63
+IV-B Results and Discussion
+Prediction Performance.
+Table II shows the classification accuracy of the compared algorithms.
+As we can see, our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert not only can provide explainability but also can achieve comparable prediction accuracy as compared to existing predictive methods.
+In addition, Lamole models show superior performance over ChemBERTa.
+This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the chemical semantics and structure-property relationships.
+Explanation Performance.
+Table III presents the explanation accuracy of the compared explainability techniques.
+It should be noted that the ground truth annotations used in our work provide additional supervisory signals.
+Therefore, we also align the generated explanations of these baselines with the annotations for fair comparison, and the ground truth annotation rate is 10%.
+Lamole improves the explanation accuracy by $1.4\%\sim 14.3\%$ compared to the baseline methods.
+The results of the explanation accuracy of Lamole under different annotation rates (10%, 20%, 50%, and 100%) on the PTC-FR and PTC-MM datasets are shown in Fig.
+5 and Fig.
+6 (a), (b).
+It is clear that more annotations can constantly enhance the accuracy of the explanation.
+Compared to Lamole ($\mathcal{L}_{M}$), only using 10% molecules with ground truth annotations can significantly improve explanation accuracy by up to 5%.
+Using more annotations (from 10% to 20%) can achieve significant improvement in explanation accuracy.
+However, raising the rate from 50% to 100% can bring a limited increase in explanation accuracy on the four PTC datasets.
+Therefore, there is a trade-off between the explanation accuracy and additional annotation costs.
+The fidelity results across the seven datasets presented in Table IV confirm that our method, Lamole, generates explanations that are more faithful to the underlying model prediction than all baselines.
+The best performing Lamole+Bert, consistently achieved the highest scores on nearly all benchmarks, including 0.61 on Mutag, 0.62 on Mutagen, and 0.68 on PTC-FM.
+This performance is substantially higher compared to dedicated explainers like OrphicX and gradient methods.
+The data indicate that Lamole’s integration of attention and gradient signals, combined with its functional-group focus, produces explanations where the removal of the identified fragments causes a maximal reduction in the model’s prediction score.
+This confirms the reliability of our approach.
+(a) Mutag
+(b) PTC-FM
+(c) PTC-MR
+Figure 5: The explanation accuracy of Lamole with different annotation rates on the Mutag, PTC-FM, and PTC-MR datasets.
+(a) PTC-FR
+(b) PTC-$M$
+(c) Molecule (ID: 155)
+Figure 6: The explanation accuracy of Lamole with different annotation rates on the PTC-FR and PTC-MM datasets, and the explanation visualization of one molecule from the Mutag dataset.
+In addition, we discuss the rationale of using labeled annotations.
+Lamole requires human-labeled annotations.
+Due to the huge knowledge base of LLMs, we explored the use of LLMs, including ChatGPT and ChemLLM [65], to annotate the ground truth.
+We input molecules’ SMILES strings into the two LLMs to ask the ground truth.
+The explanation accuracy results are shown in Table V.
+TABLE V: The explanation accuracy results when using different annotation methods
+Methods Mutag PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
+Lamole (ChatGPT) 67.6 64.5 55.0 51.0 65.6 72.7
+Lamole (ChemLLM) 62.5 72.1 61.5 57.1 71.9 72.7
+Lamole (Human) 77.8 81.1 72.2 72.0 73.1 77.3
+It is obvious that there is a significant decrease in explanation accuracy, indicating that existing LLMs may make incorrect annotations.
+Therefore, we argue that this human-in-the-loop strategy — providing slight human annotations — to guide learning is reasonable and necessary for critical scientific domains.
+We selected some representative molecules for explanation visualization.
+These explanations are shown in Figs.
+1, 6 (c), 7, and Fig.
+S-1 (see supplemental material), respectively.
+The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments.
+Compared to baseline methods, Lamole provides chemically meaningful explanations.
+Particularly, the interaction among the functional groups is successfully captured.
+As shown in Figs.
+7 (b) and S-1 (a), Lamole accurately and confidently identify benzene with amido group and benzene-1 with nitro-1 group as explanations, respectively.
+While other methods can neither provide chemically meaningful explanations nor reflect the functional group interactions.
+These explanations demonstrate Lamole’s superior interpretation in faithfully revealing the structure-property relationships.
+However, as depicted in Fig.
+S-1 (b), although the ground truth substructures, i.e., benzene-1 and amido ground, are identified, other functional groups/fragments such as carbonyl-0, $\mathrm{Br}$-0, and $\mathrm{Br}$-1 also have relatively higher importance scores.
+This may be due to complex interactions caused by multiple functional groups.
+In future work, more strategies may need to be designed to reveal such complex functional group interactions.
+(a) Molecule (ID: 156) from the Mutag dataset
+(b) Molecule (ID: 574) from the Liver dataset
+Figure 7: Explanation visualization molecules from the Mutag and the Liver datasets.
+In addition, we evaluated the performance of the compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$.
+The statistical results of $\mathrm{EP}$ on the PTC-FR, PTC-MM, and Liver datasets are presented in Figs.
+8 and S-2.
+From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property.
+Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$.
+The explanations’ plausibility $\mathrm{EP}(g)$ is defined as the ratio of the difference between the mean importance scores of ground truth and the mean importance scores of non-ground truth to the mean importance scores of non-ground truth.
+A larger ratio indicates that ground truth’s importance scores exceed non-ground truth’s.
+The high $\mathrm{EP}(g)$ values of Lamole indicate Lamole can improve the importance scores of ground truth and suppress the importance scores of non-ground truth, leading to higher confidence in matching the ground truth.
 (a) Mutag
@@
 Figure 8: The explanation plausibility of the compared algorithms on the Mutag, PTC-FM, and PTC-MR datasets.
-We selected some representative molecules for explanation visualization.
-These explanations are shown in Figs.
-1 and 4, respectively.
-The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments.
-Compared to baseline methods, Lamole provides chemically meaningful explanations.
-Particularly, the interaction among the functional groups is successfully captured.
-More visualization results can be found in Appendix.
-In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$.
-The statistical results of $\mathrm{EP}$ are presented in Figs.
-5 and LABEL:epfr.
-From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property.
-Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$.
-More analysis regarding the explanation plausibility can be found in Appendix.
-4.3 Ablation Studies
-We conducted ablation studies for each component in our Lamole.
-Specifically, we removed the attention weights in the explanations, removed the marginal loss, and only used attention weights as explanations.
-The corresponding ablation algorithms are named Lamole ($-\alpha$), Lamole ($-\mathcal{L}_{M}$), and Lamole (Att), respectively.
-The results are shown in Fig.
-6.
-The explanation accuracy decreases by 1.4$\sim$2.3when removing the attention weights.
-Removing the marginal loss can decrease the explanation accuracy by 1.0$\sim$5.0%.
-Regarding the results of using only attention weights, The explanation accuracy decreases by 1.4$\sim$6.3%.
-The above results confirm the effectiveness of using the marginal loss, attention weights, and gradients.
-Figure 6: The explanation accuracy of Lamole+Bert, Lamole+Bert ($-\alpha$), Lamole+Bert ($-\mathcal{L}_{M}$), and Lamole (Att).
-From the perspective of model training, the marginal loss enables the model to be trained under the causal signals.
-To verify, we compared the classification performance with and without the marginal loss, as shown in Table 3.
-Without the marginal loss, the classification accuracy degrades by 0.7$\sim$3.7%.
-The above results indicate that marginal loss could help identify the causal features, thereby improving classification accuracy.
-Table 3: The classification performance with and without the marginal loss (%)
-Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
+IV-C Applicability to Broader Properties
+The above experiments focus on toxicity-related properties where some ground-truth concepts are well-understood.
+To further investigate whether the explainable prediction ability applies to broader properties, we have performed additional experiments on two critical ADME properties: solubility and permeability.
+Since large-scale human-annotated datasets for ADME properties are unavailable, we employ an automated, descriptor-driven approach to generate the ground-truth functional group labels required for evaluating explanation accuracy on the ZINC dataset.
+We leverage established physicochemical principles quantified by the RDKit cheminformatics package [28] to define chemically meaningful key fragments.
+For both solubility and permeability, the process hinges on quantifying local contributions to defining global descriptors, such as topological Polar Surface Area (tPSA) for polarity and LogP for hydrophobicity.
+We utilize RDKit to calculate the contribution of each atom to the relevant property descriptor.
+We determine the ground truth structures by aggregating these atomic contributions: the contributions of all atoms belonging to a specific Group SELFIES fragment are summed.
+The fragments whose summed contributions are among the highest are then designated as the ground truth structures.
+Fragments with the highest tPSA are labeled as ground truth structures  [^1] It should be noted that the purpose of using this RDKit-based annotation is to provide computational references, rather than the true chemical causality — this needs to be rigorously verified through wet experiments..
+The results presented in Table VI and Table VII demonstrate that the advantages of the Lamole framework extend to these more diverse and complex pharmacological tasks, confirming its applicability beyond toxicity prediction.
+Regarding predictive performance (Table VI), Lamole shows substantial superiority over GNN baselines in terms of Root Mean Square Error (RMSE).
+For instance, Lamole+Bert achieved an RMSE of 0.65 for solubility.
+A similar trend holds for permeability, where Lamole+Bert achieved an RMSE of 2.1, significantly outperforming GNN baselines.
+This predictive accuracy confirms that Lamole’s architecture captures the subtle structure-activity relationships crucial for accurate ADME prediction.
+The results for explanation quality (Table VII) further show the method’s generalizability by yielding higher explanation accuracy and fidelity scores.
+The framework achieved the highest explanation accuracy (Exp.
+Acc.) for both solubility (89.1%) and permeability (73.5%), indicating that Lamole is superior at identifying the core functional groups that govern physicochemical behavior.
+Furthermore, the fidelity scores—which measure how faithful the explanation is to the model’s prediction—are also highest for Lamole variants.
+The explanation visualization shown in Fig.
+S-3 and the explanation plausibility shown in Fig.
+S-4 also confirmed the accuracy of the derived explanations.
+TABLE VI: Comparison of Prediction Performance (RMSE) on Solubility and Permeability
+Methods Solubility Permeability
+GCN [23] 0.88 3.5
+DGCNN [66] 0.85 3.4
+edGNN [21] 0.82 3.3
+GIN [62] 0.79 3.1
+RW-GNN [36] 0.77 3.0
+DropGNN [37] 0.75 2.9
+IEGN [34] 0.73 2.9
+Lamole+DistilBert 0.70 2.8
+Lamole+DeBerta 0.74 2.5
+Lamole+Bert 0.65 2.1
+TABLE VII: Comparison of Explanation Quality on Solubility and Permeability
+Methods Solubility Permeability
+Exp. Acc. ($\%$) Fidelity Exp. Acc. ($\%$) Fidelity
+SmoothGrad+GCN [48] 78.5 0.35 64.1 12.3
+CAM+GCN [67] 79.1 0.38 65.0 13.5
+GradInput+GCN [46] 77.8 0.32 63.8 11.9
+GradCAM+GCN [44] 80.3 0.41 66.2 14.0
+SmoothGrad+Bert [48] 81.5 0.43 67.5 14.5
+CAM+Bert [67] 82.2 0.45 68.1 14.8
+GradInput+Bert [46] 81.0 0.42 66.9 14.1
+GradCAM+Bert [44] 83.0 0.46 69.5 15.1
+OrphicX [30] 84.5 0.48 70.0 15.5
+GNNExplainer [64] 85.1 0.49 70.8 15.8
+PGExplainer [33] 83.9 0.47 69.8 15.3
+Gem [29] 85.5 0.49 71.0 16.0
+Lamole+DistilBert 88.4 0.52 72.1 16.5
+Lamole+DeBerta 88.7 0.50 73.2 15.9
+Lamole+Bert 89.1 0.56 73.5 17.6
+IV-D Ablation Studies
+[…]
+Figure 9: The explanation accuracy of Lamole+Bert, Lamole+Bert ($\alpha$), Lamole+Bert ($\mathcal{L}_{M}$), and Lamole (Att).
+[…]
+TABLE VIII: The classification performance with and without the marginal loss (%)
+Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
 Lamole-$\mathcal{L}_{M}$ 84.9 73.8 58.6 69.3 59.7 62.3 44.6
@@
 To investigate the attention weights, the attention weights of two molecules are depicted in Fig.
-7.
+10.
 From Fig.
-7 middle panel, it can be found the correlation among the ground truth substructures is higher than others, showcasing the rationality of using attention weights to capture the functional group interactions.
-When we aggregate the attention weights for each token, as shown in Fig.
-7 right panel, the top two attention weights of the molecule (ID:155) match the two ground truth substructures.
-However, the top two attention weights of the molecule (ID:156) do not match the two ground truth substructures.
-The above results indicate that attention weights can capture the interactions and also confirm that "attention is not explanation" (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020).
-Due to the page limitation, the limitation of the proposed work is provided in Appendix.
+10 right panel, it can be found that the correlation among the ground truth substructures is higher than others, showcasing the rationality of using attention weights to capture the functional group interactions.
+When we aggregate the attention weights for each token, the top two attention weights of the molecule (ID:155) match the two ground truth substructures (see Fig.
+10 (a) left bottom panel).
+However, the top two attention weights of the molecule (ID:156) do not match the two ground truth substructures (see Fig.
+10 (b) left bottom panel).
+The above results indicate that attention weights can capture the interactions and also confirm that "attention is not explanation" [20, 45, 1].
 Figure 10: Attention weights visualization. The ground truth substructures are highlighted in red.
-5 Conclusions
-This work proposed Lamole for explainable molecular property prediction based on language models.
-Lamole uses Group SELFIES as input for chemically meaningful semantics.
-By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations to quantify each chemically meaningful substructure’s impact on the model’s output.
-Furthermore, one marginal loss is designed to calibrate the explanations to be more faithful by aligning them with the chemists’ annotation.
-Lamole ’s effectiveness has been demonstrated through theoretical analysis and extensive experimental validation.
+Figure 11: The proposed interpretable molecular optimization pipeline integrates Lamole’s explainable molecular property prediction with an evolutionary algorithm to guide the editing process: The process begins with Lamole identifying the key functional groups in seed molecules.
+In each generation, crossover swaps the key fragments between parents, while mutation introduces novel fragments either randomly or by using a fragment vocabulary.
+The selection process retains the promising molecules for the next generation.
+In this loop, Lamole identifies new key groups and dynamically updates the fragment vocabulary for guiding the optimization.
+IV-E Computational Cost
+The pre-training stage took 11.8 hours.
+After pre-training, we fine-tuned the model on the used dataset.
+On the Mutag dataset, this process took 158s, and the evaluation time was 15s.
+For the baseline methods, SmoothGrad+GCN and OrphicX, the total training and evaluation time is 87s and 122s, respectively.
+Considering that the language model only requires pretraining once, the proposed method consumes an acceptable additional computational cost but brings comparable classification accuracy and explainability.
+IV-F Case Study: Actionable Utility for Molecular Editing
+The above experiments demonstrate that our proposed Lamole could provide chemical concepts-aligned explanations.
+This feature is crucial, particularly for practical use like molecular editing, as a concept-aligned explanation is inherently actionable for humans — these explanations can directly guide the modification of molecules.
+In this section, we illustrate the actionable utility of Lamole’s explanations in molecular editing tasks.
+Specifically, this study focuses on optimizing ADME properties.
+We achieve this by establishing an interpretable molecular optimization pipeline that integrates Lamole’s explainable molecular property prediction with an evolutionary algorithm to guide the editing process.
+The interpretable molecular optimization pipeline begins by defining the optimization objective.
+This objective corresponds to the property score, $F(\cdot)$, which serves as the fitness function for the evolutionary algorithm.
+For this study, we chose solubility and permeability as the optimization objectives.
+For instance, maximizing the solubility score or the permeability score.
+As presented in Fig.
+11, firstly, we initialize the evolutionary algorithm by selecting $N=20$ seed molecules from the ZINC dataset to form the initial population, $P_{0}$.
+For each molecule $M\in P_{0}$, the Lamole framework is immediately applied to identify the most influential functional groups that contribute to the property score, thereby informing the subsequent genetic operators.
+In each generation $t$, a new population $P_{t+1}$ is derived from the current population $P_{t}$ through the fragment-informed genetic operators.
+These operators leverage the key fragments provided by Lamole to bias the search towards chemically meaningful modifications, ensuring the evolution is both effective and interpretable.
+  • Crossover: the crossover operator is designed to facilitate the targeted exchange of critical structural information.
+Two parent molecules, $M_{1}$ and $M_{2}$, are selected from $P_{t}$.
+Instead of performing random recombination, the operator utilizes Lamole to identify the key functional groups contributing to the target property.
+These highly influential substructures are then swapped between $M_{1}$ and $M_{2}$ to generate two novel offspring, $M_{1}^{\prime}$ and $M_{2}^{\prime}$.
+This mechanism ensures that the recombination step targets the most predictive parts of the molecular structure, thereby enhancing the probability of generating molecules with desired properties in the subsequent generation.
+  • The mutation operator introduces structural diversity to explore the chemical landscape.
+We employ two distinct mutation strategies: 1) Random structural mutation: This mode ensures broad exploration.
+A molecule $M\in P_{t}$ is selected, and one of its constituent fragments encoded via Group SELFIES is arbitrarily replaced.
+This is essential for preventing the algorithm from prematurely converging to local optima.
+2) Vocabulary-based mutation: This strategy is designed to enhance the chemical plausibility and efficiency of the search by leveraging historical knowledge.
+We first define the fragment vocabulary ($V$) as the set of chemical fragments that have been identified as key contributors by Lamole.
+The vocabulary is maintained and updated throughout the evolutionary history.
+During mutation, a molecule $M$ is selected, and one of its existing fragments is substituted with a fragment retrieved from this vocabulary $V$.
+The merits of retrieving fragments from the vocabulary $V$ are twofold: (1) it ensures that the introduced structural change is biased towards known chemical building blocks that have been validated during the optimization process; (2) it balances exploration (introducing new fragments) with exploitation (recombining chemically validated motifs).
+Following the execution of crossover and mutation operators, a large pool of candidate offspring molecules is generated.
+These generated offspring are then subjected to property score evaluation.
+The target property score, $F(M^{\prime})$, for each new molecule $M^{\prime}\in P_{t+1}$ is calculated using RDKit [^2] Because we use the RDKit-computed objective, this molecular editing case is presented only as a computational demonstration, and the resulting optimized molecules do not imply real-world chemical optimality..
+Simultaneously, the Lamole model is executed on these new molecules to identify the emergent key fragments that contribute to their scores.
+If a new molecule exhibits a higher property score than its parents and contains a fragment not yet present in the vocabulary ($V$), then the fragment is updated into $V$ for use in future mutation steps.
+Finally, based on the property score $F(M)$, a selection scheme is applied to $P_{t}\cup P_{t+1}$ to determine which molecules form the population $P_{t+1}$ for the next generation.
+This entire iterative process continues until a predefined maximum number of generations $t_{max}=50$ is reached.
+To benchmark the efficacy of the proposed interpretable molecular editing pipeline, we compare our pipeline against several baselines, and all baselines are from [32].
+1) Random: This baseline applies a random noise as the perturbation to the latent representation, testing optimization performance achieved through purely random exploration in the latent space.
+2) PCA (Principal Component Analysis): This method guides modifications using the eigenvectors obtained from the PCA decomposition of the latent representation.
+This approach tests optimization guided by the directions of maximum global variance in the representation space.
+3) High-Variance: This technique attempts to target specific semantic directions by focusing on the latent representation dimension with the highest variance.
+4) MoleculeSTM: This method utilizes a multimodal structure–text model by jointly learning molecules’ chemical structures and textual descriptions via a contrastive learning strategy.
+5) Genetic Search: It directly applies genetic operators to the molecular graph representation.
+TABLE IX: Results on four molecule editing tasks, and the evaluation is the hit ratio of the property change.
+Target Property Random PCA High Variance GS-Mutate MoleculeSTM Ours
+Higher Permeability 25.23 21.36 21.98 22.00 31.76 39.55
+Lower Permeability 16.79 15.48 14.66 28.83 37.35 45.90
+More Soluble in Water 35.33 33.80 33.52 52.00 67.86 70.83
+Less Soluble in Water 43.36 39.36 42.89 47.50 64.79 66.33
+Figure 12: Visualization of the molecule editing tasks. The blue region corresponds to the edited substructures in the molecule.
+The results on the molecular editing task are summarized in Table IX, which reports the hit ratio of achieving the target property change for four distinct optimization goals.
+Across all four objectives—increasing or decreasing permeability and increasing or decreasing water solubility—our Lamole-guided evolutionary algorithm achieved the highest hit ratios.
+For enhancing permeability, our method yielded a hit ratio of $39.55\%$, surpassing the best baseline, MoleculeSTM.
+The most substantial gain was observed when optimizing for lower permeability, where our method reached $45.90\%$, higher than MoleculeSTM ($37.35\%$) and GS-Mutate ($28.83\%$).
+Similarly, for solubility, our approach demonstrated superior efficiency.
+For achieving more soluble in water, the hit ratio was $70.83\%$, narrowly exceeding MoleculeSTM and showing a wide margin over GS-Mutate.
+When optimizing for less soluble in water, our method maintained the lead with a hit ratio of $66.33\%$.
+Fig.
+12 presents four representative examples of molecular editing, illustrating the resulting structures optimized for higher solubility in water and higher permeability, respectively.
+In both cases, the properties show a notable increase following the targeted fragment modifications.
+The corresponding visualizations for edits targeting lower solubility in water and lower permeability are provided in the supplementary material.
+The performance across all objectives demonstrates that utilizing the chemically informed key fragments identified by Lamole effectively guides the evolutionary search.
+V Conclusions
+In this work, we presented Lamole, a novel framework for explainable molecular property prediction based on LMs, specifically designed to address the critical gaps in generating chemically meaningful and faithful explanations.
+To overcome the representational limitations of existing methods, we first leveraged the Group SELFIES string representation, which intrinsically encodes molecular information at the functional group level, aligning model input directly with chemical intuition.
+Furthermore, we analyzed the deficiencies of relying solely on attention or gradients and proposed an information flow-based explanation method by carefully disentangling the internal architecture of the Transformer.
+This unique approach integrates both self-attention weights and gradients to accurately quantify the contribution of each functional substructure and capture inter-group interactions.
+To ensure the plausibility of the explanation, we introduced a novel marginal loss designed to calibrate the explanations against expert annotations.
+We theoretically demonstrated that this loss aligns the generated explanations with the tangent space of the data manifold, ensuring concept-aligned and robust results.
+Empirical validation across eight diverse datasets, including toxicity, solubility, and permeability, confirmed that Lamole achieves predictive accuracy comparable to state-of-the-art baselines while significantly boosting explanation accuracy.
+We also showcased the practical, actionable utility of our framework by integrating Lamole with an evolutionary algorithm, establishing an interpretable optimization pipeline that serves as a rational guide for molecular editing and design.
+This work establishes a concept-aligned explainability framework on standard molecular benchmarks.
+However, extending it to broader generalization scenarios, including unseen chemotypes, remains an important direction for future work.